### v2

# The Hitchhiker’s Guide to Program Analysis, Part II: Deep Thoughts by LLMs

Abstract.
Static analysis is a cornerstone for software vulnerability detection, yet it often struggles with the classic precision-scalability trade-off. In practice, such tools often produce high false positive rates, particularly in large codebases like the Linux kernel. This imprecision can arise from simplified vulnerability modeling and over-approximation of path and data constraints. While Large Language Models (LLMs) show promise in code understanding, their naive application to program analysis yields unreliable results due to inherent reasoning limitations. We introduce BugLens, a post-refinement framework that significantly improves static analysis precision. BugLens guides an LLM to follow traditional analysis steps by assessing buggy code patterns for security impact and validating the constraints associated with static warnings. Evaluated on real-world Linux kernel bugs, BugLens raises precision from 0.10 (raw) and 0.50 (semi-automated refinement) to 0.72, substantially reducing false positives and revealing four previously unreported vulnerabilities. Our results suggest that a structured LLM-based workflow can meaningfully enhance the effectiveness of static analysis tools.
conference: ; Apr 2025; Earth
1. Introduction
Static analysis has long served as a cornerstone technique for identifying software vulnerabilities. By analyzing the code without dynamic execution, these techniques aim to detect various security weaknesses, such as buffer overflows and information leaks. However, static analysis tools often struggle to balance the trade-off between precision and scalability (gosain_static_2015, ; DBLP:journals/csur/ParkLR22, ).
More precise analysis, for example, symbolic execution (horvath_scaling_2024, ), can be computationally expensive and often infeasible for large codebases such as the Linux kernel (ubitect, ). Conversely, more scalable techniques sacrifice the precision for scalability, leading to a high number of false positives. For example, Suture (zhang_statically_2021, ), an advanced taint bug detection in the Android kernel shows a 90% raw false positive rate, which requires manual inspection of the results.
Specifically, its imprecision stems from two main issues:
  • $\bullet$
    Simplified Vulnerability Modeling. Static analyzers may often rely on heuristic simplified rules for vulnerability detection. For example, a static analyzer may flag every arithmetic operation as potentially overflowing. While this simplication might ensure no genuine vulnerabilities are missed, it inflates the number of false positives.
  • $\bullet$
    Over-Approximation of Path and Data Constraints. In order to avoid exponential path exploration, static analyzers often make coarse assumptions about whether a path is feasible or how data flows through the program. This over-approximation ensures analysis completes in a reasonable time, but it also flags numerous infeasible paths as potentially vulnerable, resulting in excessive false positives.
Recent advances in Large Language Models (LLMs) offer a promising avenue for overcoming these issues. Trained on vast amounts of code and natural language, LLMs exhibit remarkable capabilities in understanding code semantics, API usage patterns, and common vulnerability types (khare_understanding_2024, ; stoica_if_2024, ; fang_large_2024, ). By leveraging these broader insights, LLM-based approach might able to: (1) enhance vulnerability modeling by providing a more nuanced understanding of code semantics, and (2) refine path and data constraints by a selective analysis of semantically plausible paths and data flows.
However, LLMs are not a silver bullet for program analysis. Despite their semantic understanding capabilities, LLMs are not inherently equipped for the rigorous demands of program analysis (qian_lamd_2025, ; he_code_2024, ). Their reasoning proves brittle, particularly when confronted with the complex program dependencies crucial for security analysis (kambhampati_can_2024, ; chen2025reasoning, ; chapman_interleaving_2025, ). Indeed, our initial experiments confirm that naively applying LLMs to program analysis, for instance, by simply asking “Is this warning a true positive?”, yields highly unreliable results, frequently misclassifying vulnerabilities or failing to identify critical flaws. This is often because LLMs tend to fixate on surface-level code features, missing the critical dependencies that dictate program behavior and security properties, especially within intricate control and data flows.
Our work addresses this challenge by introducing a structured guidance framework that directs LLM reasoning according to static program analysis paradigm. This approach rests on the premise that static analysis, while theoretically sound, is limited by practical tradeoffs in precision and scalability. By situating LLM reasoning within this established methodology, we “compel” the model to analyze more rigorously than it would by default, thus mitigating the inherent limitations of LLMs in code reasoning.
In this work, we introduce BugLens, an innovative framework that post-refines the results of static analysis using LLMs. Rather than blindly applying LLMs to analyze program, BugLens is carefully orchestrated to teach LLMs key concepts of program analysis and guide them toward reasonable analytical procedures. By analyzing the output of static analysis, BugLens  complements the limitations of existing tools, especially in terms of precision, and yields more accurate and actionable vulnerability detection for practical codebases. We demonstrate that this combined approach significantly improves the precision of taint-bug detection in the Linux kernel, reducing the need for manual inspection of false positives, and even uncovering previously ignored vulnerabilities.
We summarize our contributions as follows:
  • $\bullet$
    Post-Refinement Framework. We introduce BugLens, a post-refinement framework that supplements traditional static analysis to boost precision of the results, overcoming various practical weaknesses identified from real-world static analysis tools.
  • $\bullet$
    Structured Analysis Guidance (SAG). We design a structured workflow that directs LLMs to follow the principles and processes of established static analysis paradigms. This guided approach achieves better vulnerability detection compared to naive LLM prompting strategies.
  • $\bullet$
    Empirical Results. Evaluated on the Linux kernel, our solution improved the precision of a real-world static analysis tool from 0.1 to 0.72. Interestingly, prior manual analysis incorrectly filtered four warnings that were retained by BugLens.
2. Background & Motivation
2.1. Taint Bugs & Static Analysis
A taint bug is a type of vulnerability that arises when untrusted data is improperly handled within a program, leading to potential security risks. For example, in the Linux kernel, the ioctl system call often handles user input, without proper validation vulnerabilities can occur, such as buffer overflows and infinite loops (machiry_dr_2017, ; zhang_statically_2021, ).
We focus on taint bugs for several reasons. First, a wide variety of security vulnerabilities can be modeled as taint problems, including buffer overflows, integer overflows, use-after-free, and other memory corruption issues. Second, taint analysis provides a well-defined conceptual framework for reasoning about data flow from untrusted sources to sensitive operations (sinks). Third, the taint static analysis is both practical and widely adopted in security tools. These make taint analysis an ideal foundation for our work.
Suture (zhang_statically_2021, ) is a state-of-the-art static analysis tool designed specifically for detecting taint-style vulnerabilities in the Linux kernel. It employs a taint-tracking approach to identify potential vulnerabilities by analyzing the flow of untrusted data through the kernel code. Suture precisely tracks how data from user inputs, such as those passed via the ioctl system call, propagates through various functions and data structures, flagging any instances where this tainted data could lead to security issues based on their detectors.
Notably, Suture is an advanced taint analysis tool that can even analyze multi-entry data flows, which means it can track tainted data propagating through multiple system calls. Nevertheless, it still faces challenges in accurately determining whether the identified potential bugs are indeed exploitable, and still require considerable manual review to confirm their validity. We discuss these challenges in detail in the following parts of this section.
Figure 1. A potentially vulnerable code snippet from the Android kernel, simplified for illustration. The array values is tainted and all its elements are tainted as well.
[…]
This code snippet is a simplified version of a function from the Android kernel. It shows how the function retrieves the number of channels and their types, which could be exploited if not properly validated.
2.2. Motivating Example
The primary motivation for BugLens stems from the persistent challenge of achieving both precision and scalability in static vulnerability analysis, particularly for large-scale, complex systems like the Linux kernel. While static analysis is indispensable for identifying potential security issues, current tools often generate a high volume of findings, many of which are false positives (FPs) upon manual inspection. This significantly burdens security analysts and developers. BugLens aims to bridge this gap by introducing a methodology for more precise, context-aware evaluation of potential taint-style vulnerabilities, leveraging Large Language Models (LLMs) in a guided manner.
Figure 1 shows a simplified code snippet from the Android kernel. The function is called with an array of tainted variables *values (Note: the address represented by values is not tainted, but its contents are, from a copy_from_user). Thus, retrieved by GET_NEXT, the variable nums and command_id are tainted as well. Static analysis may flag potential bugs for the following detectors:
A standard static taint analyzer might flag the following issues:
  • $\bullet$
    Tainted Loop Bound (TLB): At line 10, the loop iterates based on the tainted variable nums. A very large value for nums (e.g., close to INT_MAX) could lead to excessive iteration, potentially causing a Denial-of-Service (DoS) or creating conditions for other vulnerabilities.
  • $\bullet$
    Tainted Pointer Dereference (TPD): At lines 11, the GET_NEXT macro is called. Inside this macro (line 4), the tainted pointer values (represented as ptr) is dereferenced (*ptr++) after being incremented. Since the number of increments depends on the loop controlled by tainted nums, the pointer values in the iteration is tainted as well (an implicit taint propagation). Therefore, it could be reported as a tainted pointer dereference, which may potentially lead to an out-of-bounds or arbitrary memory access.
2.2.1. Ground Truth Analysis
However, a deeper analysis reveals a more nuanced reality:
  • $\bullet$
    The TLB is a potential true positive. If nums is extremely large, but the data in values remains within the bounds defined by MAX_PP_PARAMS_SZ for many iterations, the loop can run for a long time. If the read command_ids are invalid (triggering the default case and continue on line 18), the if (rc) check on line 20 is bypassed for those iterations. The loop only terminates when i reaches nums. An extremely large nums could thus cause excessive computation, leading to a DoS.
  • $\bullet$
    The TPD is a false positive. The potential out-of-bounds read is prevented by the explicit bounds check within the GET_NEXT macro itself (line 3). It compares the incremented pointer address against the calculated param_max_offset (derived from the known buffer size MAX_PP_PARAMS_SZ) before the dereference occurs (line 4). If the check fails, rc is set, and the dereference is skipped.
2.3. Challenges in Static Analysis
Our motivating example, involving the analysis of TLB and TPD (illustrated in Figure 1), highlights the specific challenges stemming from this precision gap, which BugLens aims to address.
2.3.1. Challenge 1: Simplified Vulnerability Modeling
A primary source of imprecision is the reliance on simplified vulnerability detection modeling. To maintain scalability and avoid missing potential bugs, static analyzers may use overly simplified rules. For instance, an analyzer might flag any tainted pointer in the loop bound conditions as a potential DoS vulnerability, without considering the properties of the loop, such as the TLB example. Similarly, flagging all arithmetic operations involving tainted variables as potentially overflowing ensures coverage but generates noise.
The difficulty lies in defining and implementing effective and comprehensive rules within traditional static analysis frameworks. Vulnerabilities manifest in diverse and subtle ways. Consider the TLB example: a tainted loop bound might lead to Denial-of-Service (DoS) directly through excessive iterations, e.g., while (i < nums) where nums is tainted and set to a large value. Alternatively, the DoS might also happen when the tainted data is used in loop control statement such as continue or break. Precisely encoding rules to capture the vast number of unsafe code patterns, subtle contextual factors, and the impact of complex control flow is extremely challenging. As a result, this reliance on simplified heuristics inevitably leads to false positives in static analysis for security vulnerability detection.
2.3.2. Challenge 2: Over-Approximation of Constraints
The second major source of imprecision is the over-approximation of path and data constraints. To avoid the exponential complexity of exploring all possible execution paths and data states, static analyzers make coarse assumptions about path feasibility and data flow. While essential for scalability, this over-approximation means the analysis often considers execution paths that are actually infeasible in practice.
Consequently, warnings are generated for potential vulnerabilities on these infeasible paths, significantly inflating the false positive count. For example, an analyzer might flag the TPD scenario as dangerous without determining if the tainted pointer has been properly checked with its legal bounds before dereferencing. This imprecision results in a high volume of false positives, and forces developers to manually trace complex paths and data constraints to determine feasibility, undermining the efficiency of automated analysis.
2.4. The Opportunity in LLMs
Recent advances in Large Language Models (LLMs) offer a promising avenue to address these challenges, particularly in post-refining the results generated by scalable static analyzers. Trained on vast amounts of code and natural language, LLMs exhibit remarkable capabilities relevant to overcoming the limitations described above:
  • $\bullet$
    Enhancing Vulnerability Modeling (Addressing C1): LLMs possess a nuanced understanding of code semantics, common programming patterns, and API usage (li2025codei, ; ni2024next, ; liu_generating_2025, ; khare_understanding_2024, ). They can potentially recognize complex or subtle vulnerability indicators that are difficult to capture with predefined simplified heuristics. By analyzing code context (including comments and variable names), LLMs might differentiate between genuinely unsafe patterns and benign ones.
  • $\bullet$
    Refining Path and Data Constraints (Addressing C2): While not performing formal reasoning, LLMs can leverage their semantic understanding to assess the feasibility of paths and data flows flagged by static analysis. They might leverage contextual information to determine if the flagged bug is likely infeasible in practice, thereby helping to prune false positives arising from coarse path and data constraint analysis.
2.4.1. Challenge 3: Reasoning Hurdles for LLMs.
Despite their potential, LLMs are not inherently suited for rigorous program analysis tasks out-of-the-box. Their reasoning can be fragile when dealing with complex program dependencies, and they lack systematic constraint-solving mechanisms. Simply asking an LLM a high-level question like “Is this static analysis warning a true positive?” often yields unreliable results.
Specifically, LLMs (even the most advanced ones, see §5.4) may struggle with intricate control and data flow dependencies. For instance, when analyzing the TLB example (Figure 1), an LLM might observe sanity checks on the tainted data and incorrectly conclude the warning is a false positive:
”There are sanity checks for the bound value… if the check fails, an error code is set, and the function jumps to cleanup… Therefore, the tool’s warning appears to be a false positive.”
However, as shown in the example, the LLM might fail to recognize that a continue statement within the loop body bypasses these checks under certain conditions, potentially allowing the infinite loop to occur. This highlights a critical limitation: LLMs may overlook the precise implications of control flow structures and data dependencies, leading to flawed conclusions.
Therefore, effectively leveraging LLMs for refining static analysis requires more than simple prompting. It necessitates a structured approach that guides the LLM’s reasoning process, teaching it relevant program analysis concepts and focusing its attention on critical code features and dependencies. This challenge motivates the design of our BugLens framework, which aims to orchestrate LLM analysis for reliable post-refinement.
2.5. Our Approach
The challenges outlined in our motivation, namely the imprecision stemming from Simplified Vulnerability Modeling (C1) and Over-Approximation of Constraints (C2) in scalable static analysis, and Reasoning Hurdles for LLMs (C3) necessitate a new strategy to tackle these challenges. To address these, we introduce BugLens, a framework designed specifically as a post-refinement for results from traditional static analyzers.
BugLens’s primary goal is to enhance the precision of static analysis by filtering their imprecise analysis results. The framework comprises the following components, each targeting specific challenges:
  • $\bullet$
    Security Impact Assessor (SecIA): Addressing Simplified Vulnerability Models (C1). Instead of relying solely on predefined patterns, it uses LLMs to analyze the potential security impact if tainted data identified by static analysis were completely controlled by an attacker. It then evaluates whether the tainted value could potentially lead to security vulnerabilities (e.g., memory corruption, DoS), as detailed in §3.2. By focusing on semantic consequences rather than simplified heuristics, SecIA provides a more precise assessment of security impact, overcoming C1.
  • $\bullet$
    Constraint Assessor (ConA): Addressing Over-Approximation of Constraints (C2). ConA performs a targeted analysis of the (semantically) relevant constraints (e.g., sanitizers) along the data flow path (§3.3). It first identifies and collects the relevant constraints, then assesses whether these constraints effectively prevent the triggering of the specific vulnerabilities. The evaluation helps filter out false positives due to the over-approximation of path and data constraints, which are inherent in scalable static analysis (C2).
  • $\bullet$
    Structured Analysis Guidance (SAG): Addressing Reasoning Hurdles for LLMs (C3). Simply prompting LLMs for analysis is unreliable due to their potential to overlook complex control and data dependencies. SAG is designed to mitigate this (C3). It provides structured analysis guidance to the LLMs performing ConA. Leveraging domain knowledge from program analysis, SAG employs carefully constructed prompts and few-shot examples. SAG instantiates key analysis principles, demonstrating how to systematically dissect code, trace dependencies, and evaluate conditions, especially in complex scenarios. This guidance steers the LLM towards a more rigorous and reliable analysis process, making the use of LLMs within BugLens effective.
3. Design
3.1. Overview and Approach
3.1.1. Problem Scope & Goal
In this paper, we focus on identifying and analyzing taint-style vulnerabilities in operating system kernels. These vulnerabilities arise when untrusted data flows from external sources (e.g., ioctl() calls) to security-critical operations. A security-critical operation could be memory corruption such as buffer overflow, infinite loop, or other dangerous operations that could compromise the kernel when manipulated with malicious input.
As mentioned in §2, the goal of BugLens is to facilitate the (scalable but imprecise) static analysis of taint-style vulnerabilities by leveraging LLMs in a more effective way.
3.1.2. Direct Prompt & Design Rationale
Our investigation reveals fundamental limitations in simply asking LLMs to evaluate static analysis results directly. When presented with potential vulnerabilities identified by static analysis tools, a naive approach would be to directly prompt the LLM with questions like “Is this code vulnerable?” while providing the static analysis report and relevant code context. We term this basic approach the Direct Prompt.
Our experiments (detailed in §5.4) demonstrate that such Direct Prompts lead to a concerning number of false negatives — cases where the LLM incorrectly classifies actual vulnerabilities as safe code. The Direct Prompt approach, while intuitive, proves unreliable for vulnerability detection.
We hypothesize that this limitation stems from the fundamental nature of current LLMs. Trained primarily on predicting the next token based on vast datasets, these models excel at recognizing common syntactic and semantic patterns. However, this strength can become a weakness in complex analysis tasks. LLMs may develop heuristics based on surface-level features (mccoy_embers_2024, ); for example, identifying the presence of a “sanity check” construct often correlates statistically with safe code in training data. Consequently, the model might classify code containing such a check as safe without performing the deeper reasoning required to determine if the check is actually effective under all relevant execution paths. Real-world vulnerabilities often exploit exactly these scenarios: checks that are bypassable, incomplete, or rendered ineffective by intricate control and data flows. The model’s tendency to rely on learned statistical correlations (prabhakar_deciphering_2024, ), potentially driven by attention mechanisms prioritizing frequent patterns, may cause it to overlook these critical dependencies.
To address these reasoning limitations, we propose Structured Analysis Guidance (SAG) - a structured approach for guiding LLMs toward more rigorous and effective vulnerability analysis:
  • $\bullet$
    Guided Stepwise Vulnerability Analysis: We design Constraint Assessor (ConA) as a four-stage process that mirrors static analysis process. This stepwise workflow guides the LLM through: (1) identifying vulnerability reachability conditions, (2) collecting relevant constraints along taint paths, (3) summarizing each constraint’s preconditions and postconditions, and (4) evaluating whether these constraints effectively prevent the vulnerability. This systematic approach helps overcome the LLM’s tendency toward shortcut reasoning. We detail this workflow in §3.3.
  • $\bullet$
    Guided Path Condition and Data Constraint Analysis: To address challenging analysis where implicit path conditions and data constraints often elude naive prompting, we provide the LLM with detailed guidance and few-shot examples, which explicitly direct the LLM through analyzing complex control flow constructs, distinguishing between different types of constraints, and properly evaluating their effectiveness. This in-context learning approach helps the LLM identify subtle vulnerabilities. We elaborate on these techniques in §3.4.
Figure 2. Overview of BugLens, showing (1) Security Impact Assessor (SecIA) first assesses the security impact of the potential bugs identified by static analysis, and then (2) Constraint Assessor (ConA) assesses the data constraints and evaluates if the bug is feasible. ConA is guided by (3) Structured Analysis Guidance (SAG) to reason on the code more effectively, and can interact with the (4) Project Knowledge Agent (PKA) to get information about the codebase on-demand.
[…]
Overview of our workflow, showing (1) static analysis and (2) LLM-based filtering
3.1.3. Design Components
Besides prompting LLMs with SAG, the rest of the BugLens framework consists of three components, as shown in Figure 2:
  • $\bullet$
    Security Impact Assessor (SecIA): This component evaluates the security impact of the potential bugs identified by static analysis. It identifies the Critical Operations (COp) that are influenced by the tainted data and filters out benign patterns.
  • $\bullet$
    Constraint Assessor (ConA): This component performs a multi-step analysis to evaluate the feasibility of the potential bugs. It collects the path conditions and data constraints, summarizes them, and evaluates whether they are effective in preventing the vulnerability.
  • $\bullet$
    Project Knowledge Agent (PKA): This component allows the LLM to access the codebase on-demand, enabling it to retrieve global codebase information.
Additionally, BugLens also adopts some commonly-used prompting techniques, such as (1) Majority-vote querying, where we query the model multiple times and take the most common answer; (2) Chain-of-Thought (CoT) prompting (wei_chain–thought_2023, ); and (3) Schema-constrained summarization, where a follow-up prompt requests the model’s own output in a fixed XML format, making LLM’s response easy to parse for subsequent steps.
3.2. Security Impact Assessor (SecIA)
The core insight behind SecIA is based on the fundamental evaluative question: What are the consequences if tainted data assumes arbitrary values?
3.2.1. Core Assumption and Rationale
SecIA operates on a fundamental assumption regarding attacker capability at the point of initial impact assessment: Aribitary Control Hypothesis (AC-Hypo). For a given program location $K$ where static analysis reported an operation $Op(v)$ involving tainted data $v$ (the sink), (1) we hypothesize that an attacker can control $v$ to take any value $v_{atk}$ (within the constraints of its data type), and (2) we provisionally ignore any effects of checks or path conditions (even explicit checks) encountered on analysis. In other words, we assume that the attacker can take any value to anywhere (successors of the sink node in the control flow graph).
This hypothesis enables SecIA to streamline analysis. By assuming the attacker achieves both arbitrary value control and can always reach the potential vulnerability, SecIA focuses solely on potential security impact. This permits early filtering of findings based purely on consequence, (safely) reducing subsequent analysis load. Critically, this approach defers the complex validation of actual program constraints—including path feasibility, value ranges, and importantly, whether those potentially protective checks or sanitizers can be manipulated through tainted data. This deferral mitigates false negatives (FNs) by preventing premature dismissal of vulnerabilities due to reliance on potentially bypassable checks or inaccurate LLM constraint reasoning about their effectiveness. The effectiveness of this design in reducing FNs is validated experimentally in §5.4.2.
For instance, in our motivating example, a simple analysis might deem the Tainted Loop Bound (TLB) on nums safe (i.e., not a bug) due to the goto statement potentially preventing an infinite loop condition. On the other hand, SecIA’s approach, guided by the AC-Hypo, intentionally ignores the effect of the goto statement at this stage (as its condition depends on the tainted data command_id) and proceeds to identify potential downstream effects (e.g., DoS) resulting from an arbitrarily large nums. The analysis of whether the goto check can indeed be bypassed under feasible program conditions is deferred.
3.2.2. Workflow
SecIA analyzes each location $K$ reported by the static analysis Taint Sink finding:
  • $\bullet$
    Forward Influence Analysis: Suppose the tainted data $v$ is hypothetically controlled by an attacker as $v_{atk}$, and the $Op(v)$ is a taint sink $K$ identified by static analysis. Starting from the result of $Op(v_{atk})$ at location $K$, trace how this tainted value propagates or influences subsequent program execution.
  • $\bullet$
    Identify Influenced Critical Operations: From the forward analysis, SecIA first pinpoints critical operations that are affected by the hypothetical attacker-controlled value $v_{atk}$. The analysis focuses on operations relevant to known vulnerability classes such as memory safety (e.g., arbitrary memory access) and DoS (e.g., insecure loop bounds).
  • $\bullet$
    Filter Benign Operations: SecIA leverages LLM-driven semantic understanding to distinguish and filter out operators that are actually benign. For instance, while the system flags potential risks when $v_{atk}$ impacts sensitive operations like array indexing (e.g., a[v]), it deliberately ignores cases recognized as safe, such as routine data structure traversals or searches (e.g., comparing element.id to v, a complete example appaers in §5.6). This selective filtering directs subsequent analysis (§3.3) to concentrate on issues that are more likely to present genuine security risks.
  • $\bullet$
    Output. The set of influenced Critical Operations identified in the last step. This set represents the downstream locations potentially impacted by the originally tainted data $v$ under the arbitrary control assumption.
In the Forward Influence Analysis step, the LLM is leveraged to identify potential influences, including indirect taint propagation. For example, in the motivating example, the LLM can recognize that the incremental pointer values is indirectly influenced by the tainted data nums through the GET_NEXT(). By doing so, SecIA can even identify unique bugs that are not directly related to the taint source (albeit we currently don’t have any such bugs in our experiments).
3.3. Constraint Assessor (ConA)
The Constraint Assessor (ConA) aims to identify whether the bug of COp can be triggered by analyzing the data constraints.
3.3.1. Design Rationale: Precondition and Postcondition of Constraints
As potential vulnerabilities often persist because the data constraints intended to prevent them are not truly effective in the specific context where the tainted data is used §3.1.2. Simply identifying the presence of a check or sanitization routine is insufficient. To rigorously evaluate effectiveness and guided by the paradigm of formal methods, the ConA is designed to determine:
  • $\bullet$
    Precondition. A constraint might exist in the code but be bypassed on a specific execution path leading to the potential vulnerability, or its activation might depend on configuration or state variables that are not set appropriately. Analyzing the Precondition, the conditions necessary for the constraint’s logic to execute addresses this.
  • $\bullet$
    Postcondition. Even if a constraint is activated, it might not be strict enough to prevent the specific value or range required to trigger the bug (e.g., checking x < 200 is ineffective if the vulnerability can happen if x > 100). Analyzing the Postcondition, the effect on the data’s possible values or state after the constraint operates addresses this.
Therefore, collecting and reasoning about both the Preconditions and Postconditions of constraints along the taint flow is essential for ConA to move beyond superficial checks and accurately assess whether the existing safeguards collectively guarantee the elimination of the vulnerability at the critical operation. This detailed steps is described in the following sections.
3.3.2. Overview of Workflow
As shown in Figure 3, the Constrant Assessor involves a four-step, LLM-guided workflow:
Figure 3. The workflow of Constraint Assessor (ConA)
[…]
This figure illustrates the multi-stage workflow of ConA .
  • $\bullet$
    Step 1: Critical Operation Reachability Analysis. The process begins by determining the conditions under which program execution can reach the specific Critical Operation (COp) location. This establishes the base requirements for the vulnerability to be possible.
  • $\bullet$
    Step 2: Backward Constraint Collection. Next, the analysis traces the tainted data flow path(s) backward from the COp towards the data’s source. Along this path, it identifies code segments—such as conditional statements, assertions, or calls to validation functions—that appear intended to act as constraints on the tainted data’s value or range before it is used at the COp.
  • $\bullet$
    Step 3: Constraint Effect Summarization. Each potential constraint identified in Step 2 is then analyzed in detail. This step aims to understand the constraint’s specific effect: Under which conditions (its activation precondition) does the constraint apply, and what impact does it have on the tainted variable’s possible numerical range (its range effect postcondition)?
  • $\bullet$
    Step 4: Final Vulnerability Evaluation. Finally, it performs a reasoned evaluation to determine if the vulnerability can be triggered. If still triggerable, the finding is classified as a “Potential Vulnerability.” Otherwise, if the LLM determines that the constraints effectively prevent the vulnerability from being triggered, the finding is classified as “Eliminated.”
This workflow leverages LLMs to interpret code, identify relevant patterns and constraints, and perform the heuristic reasoning required for each step. The following subsections (§3.3.3 through §3.3.6) elaborate on the specific design, rationale, heuristic considerations, and soundness implications inherent in each stage of this analysis process.
3.3.3. Step 1: Critical Operation Reachability Analysis.
The analysis begins by determining the conditions required for program execution to reach the specific Critical Operation (COp) location previously identified by SecIA as potentially vulnerability. These reachability conditions form the base constraints that must be satisfied for the vulnerability to be triggerable via this COp.
Condition Representation: When analyzing path conditions, we allow the LLM to generate natural language summaries that capture the semantic meaning of complex reachability conditions. For instance, where traditional analysis might struggle, an LLM might identify a condition like: “The operation is only executed if init_subsystem() returned zero (success), AND the device state in dev->status equals STATUS_READY.”
The LLM is then tasked (in Step 4) to evaluate these semantic constraints to assess potential impact. The natural language representation allows the LLM to leverage its understanding of the code’s intent and context, but introduces a higher degree of heuristics and potential uncertainty into the final evaluation, which must be acknowledged.
3.3.4. Step 2: Backward Constraint Collection.
Once the Critical Operation (COp) and its local reachability conditions are identified, the next step is to gather potential constraints imposed on the tainted data before it reaches the COp. This involves tracing the data flow path(s) for the tainted variable backward from the COp toward its source(s).
Leveraging Code Context and Structure: The LLM is tasked with identifying validation and sanitization operations of the tainted data. Specifically, the LLM looks for:
  • $\bullet$
    Conditional branches (if, switch) whose conditions involve the tainted variable or related data.
  • $\bullet$
    Assignments or operations that transform the tainted data.
  • $\bullet$
    Function calls that take the tainted data (or related data) as arguments.
Heuristic Identification in Domain Context: Unlike formal symbolic execution, which aims to exhaustively collect all mathematically precise path conditions, this LLM-based collection is heuristic. It may not identify every constraint on all possible paths. However, we leverage the LLM’s pattern recognition capabilities, particularly within our target domain (Linux kernel). The prompt encourages the LLM to pay attention to common kernel patterns, naming conventions (e.g., functions named check_..., validate_...), and comments that often signify sanitization or validation routines, even if they are complex to model formally. This allows the LLM to identify likely and semantically significant constraints that purely syntactic approaches might suffer from path explosion.
Example: (Figure 3) Tracing back from critical_op(x) in caller(), the LLM identifies the call check_x(...) to be a potential constraint on x. It would then query the Information Agent for the definition of check_x() to analyze its body (as detailed in Step 3)
3.3.5. Step 3: Constraint Effect Summarization
After collecting potential constraining code segments (like the function check_x) in Step 2, this step analyzes the effect of these segments on the tainted variable. The goal is to understand how different execution paths within these segments modify the possible range of the tainted variable, and under which conditions (preconditions) those paths are taken.
Traditional formal methods, like symbolic execution combined with abstract interpretation, analyze such segments rigorously. However, to ensure soundness and termination, particularly with loops or complex arithmetic, they often employ over-approximations for numerical ranges. While guaranteeing soundness (no false negatives), this necessary conservatism can sometimes limit precision, failing to prove tight constraints that actually hold, thus potentially leading to false positives later.
BugLens utilizes the LLM. We prompt the LLM to interpret the control flow and data transformations within the identified constraint segment to heuristically estimate the range effects. The prompt asks the LLM to first identify all major execution paths through the provided code segment (e.g., the body of check_x), and then:
  • $\bullet$
    For each path, determine the Path Condition (Precondition): The conditions on the parameters or relevant program state required to take the path.
  • $\bullet$
    For each path, determine the Estimated Effect on Tainted Data (Postcondition): Analyze operations like assertions (assert(x > 0)), assignments (x = sanitize(x)), and comparisons to infer the resulting constraint on the tainted variable’s range.
Table 1. The precondition and postcondition pairs for the check_x function. The oc and c stand for other_config and config respectively.
Path Precondition Postcondition
$\pi_{1}$ oc = SKIP $x\in(-\infty,+\infty)$
$\pi_{2}$ oc $\neq$ SKIP $\land$ c = 1 $x\in(0,100)$
$\pi_{3}$ oc $\neq$ SKIP $\land$ c = 2 $x\in(0,+\infty)$
$\pi_{4}$ oc $\neq$ SKIP $\land$ c $\notin\{1,2\}$ $x\in(-\infty,+\infty)$
Example: (Figure 3) For the function check_x in our example, the goal is for the LLM to output a summary similar to Table 1, which maps path conditions (preconditions) to range effects (postconditions). The path $\pi_{1}$ and $\pi_{4}$ are bypass paths, which means that the check_x function will not effectively limit the range of x.
Heuristic Nature and Soundness Trade-offs: This step relies on the LLM’s code interpretation capabilities and is heuristic, particularly in its estimation of range effects and its handling of complex control flow. Similarly to the Step 1, these summaries could also be represented in natural language, which brings more flexibility but also more uncertainty.
Rationale: We employ this LLM-based heuristic approach acknowledging the trade-offs. It aims to overcome the potential precision limitations of conservative formal abstractions and achieve broader applicability in the complex Linux kernel domain, where precise, scalable formal analysis is challenging. The goal is significantly improved precision (fewer FPs) compared to baseline static analysis, accepting a carefully managed risk of heuristic errors (potential FNs). The practical rate of False Negatives is thoroughly examined in our experimental evaluation in §5.3 and §5.5.
3.3.6. Step 4: Final Vulnerability Evaluation.
This final step synthesizes the findings from the previous analyses to determine if the identified constraints effectively neutralize the potential vulnerability associated with the Critical Operation (COp). The goal is to classify the initial finding as either “Eliminated” (constraints provably prevent the the vulnerability happening) or ”Potential Vulnerability” (no such guarantee found).
Specifically, the LLM is guided to perform the following assessment:
  • $\bullet$
    Analyze Constraint Activation: The first step is to determine if the precondition (after considering user/kernel control effects) is implied by the known COp’s Reachability Conditions. If it is, the constraint is active as an effective constraint. Otherwise, we need to further analyze its activation based on the following heuristic.
  • $\bullet$
    Heuristic to Untainted Data: For untainted data, use the heuristics (e.g., variable name, __user flag) to determine if the variable can be controlled by user, then use the worst-case assumption that the user can manipulate these variables and analyze if the precondition can be unsatisfied. And a possibility of unsatisfied precondition makes the constraint disqualified.
  • $\bullet$
    Analyze Constraint Effectiveness: Next is to determine if the postcondition of the remaining active constraints
    prevents the Bug Condition of the COp from ever being satisfied. For example, if the vulnerability happens when idx < 0 and the constraint’s effect is estimated as idx in $[0,100]$, the vulnerability is prevented.
Soundness Implications: This final evaluation relies on the LLM’s ability to perform heuristic logical reasoning. As a result, this step may introduce the following concerns: (1) The heuristic to untainted data (worst-case assumption for user control) may underestimate/overestimate constraint effectiveness, (2) errors in the LLM’s logical reasoning, range comparisons, or handling of complex conditions could lead to classification errors, and (3) it inherits any potential unsoundness from Step 3 caused by incorrect range estimations or missed bypass conditions .
3.4. Structured Analysis Guidance (SAG)
To enable the LLM to effectively perform the code analysis tasks in ConA, we employ Structured Analysis Guidance (SAG) — a prompting strategy with structured reasoning templates and few-shot examples. SAG provides: (1) Guided Stepwise Vulnerability Analysis with step-by-step instructions that decompose complex analysis tasks following established formal method principles (the analysis of precondition and postcondition, as described in §3.3), and (2) Guided Path Condition and Data Constraint Analysis demonstrating how to analyze challenging data constraints and path conditions from code, as described below.
3.4.1. Guided Path Condition Analysis
In the analysis of the path condition in Step 1 and Step 3 of ConA3.3.3 and §3.3.5), we guide the LLM using prompts designed to extract path conditions from the source code surrounding the operation of interest. For example, consider the path leading to the op(x) call within the caller function:
SAG asks the LLM to identify and categorize:
  • $\bullet$
    Bypass Conditions: Identify conditional statements where taking a specific branch avoids the operation. The LLM is instructed to extract the condition and negate it to find the requirement for not bypassing the operation.
    Example: The condition x > 100 leads to invalid, bypassing critical_op(x). The negated condition required to proceed towards the OP is x $\leq$ 100.
  • $\bullet$
    Direct Conditions: Identify conditional statements where taking a specific branch is necessary to reach the operation along the current path. The LLM extracts the condition directly.
    Example: Reaching sink(x) requires entering the if block, so the condition is other_config $\neq$ SKIP.
[…]
3.4.2. Guided Data Constraint Analysis
In the analysis of data constraints in Step 2 and Step 3 of ConA3.3.4 and §3.3.5), We consider the following data constraints:
  • $\bullet$
    Type constraints. The variable’s static type already restricts its range (e.g., uint8 is always in the range of $[0,255]$).
  • $\bullet$
    Validation (transferable to source). The program tests the value and aborts or reports an error if the test fails, without modifying the value. Because the check refers to the current value, the knowledge gained from this check (e.g., “the value must be $\geq 0$ on the success branch”) also applies to all source variables that influenced this value in the data flow.
  • $\bullet$
    Sanitization (not transferable to source). The program writes a new, corrected value back to the variable (e.g., clamping it to a range). This operation severs the connection to the original value, so any property we learn afterwards applies only to the sanitized copy, not to the original source variables in the data flow.
The key difference is that validation knowledge travels backward along the data-flow graph, while sanitization overwrites the flow and stops the transfer.
[…]
3.5. LLM Agent for Codebase Information
Following recent work (llift, ), we adopt an agentic design where the LLM acts as an interactive code analyst. It begins with a single function from the repository, identifies potential issues, and explicitly requests more context (e.g., other function definitions, struct layouts, global variables) if needed. The system then retrieves the relevant code snippet(s), and this iterative loop continues until the LLM has sufficient context to complete the analysis.
We provide a set of request types (function definitions, struct layouts, global variables, etc.), enabling the LLM to gather broader codebase information at any point. This design is both flexible and extensible: new request types can be supported by adding corresponding backend callbacks.
To facilitate these interactions, the knowledge retrieval system parses requests, locates relevant code, and formats responses for subsequent reasoning steps. We implement this system (about 500 LOC in Python) on top of CodeQuery (kopathy_ruben2020codequery_2025, ), extending its default functionality with custom handlers for specialized queries.
4. Implementation
BugLens is implemented with approximately 7k tokens (30k characters) in prompts (detailed in §B) and 2k lines of Python code that manages API requests and codebase querying functionality.
4.1. Inferring Variable Names
Static analysis tools typically operate on compiler-generated IR code (like LLVM IR), which differs significantly from source code representation. Since BugLens analyzes source code directly, we need to map variables between these representations.
The IR code provides line numbers that correspond to locations in the source code, along with data flow information. However, compiler optimizations often create, eliminate, or transform variables, making this mapping non-trivial. For example, LLVM might convert a simple ternary expression into a complex bitwise operation.
We leverage LLMs to perform this variable name inference by providing them with: (1) the line number from source code and (2) the data flow from IR code. The LLM then identifies the corresponding source-level variables, effectively bridging the gap between IR-level analysis and source code understanding.
5. Evaluation
5.1. Research Questions
Our evaluation aims to address the following research questions.
  • $\bullet$
    RQ1: (Effectiveness) How effective is BugLens in identifying vulnerabilities?
  • $\bullet$
    RQ2: (Component Contribution) How does the the prompt design affect the performance of BugLens? especially for the SecIA and SAG component?
  • $\bullet$
    RQ3: (Model Versatility) How does the performance of BugLens vary across different LLMs?
5.2. Experimental Setup
We primarily evaluate BugLens using OpenAI’s o3-mini model, specifically o3-mini-2025-01-31. This model was chosen as our primary focus because, as demonstrated in our evaluation for RQ3 (Section 5.5), it achieved the best overall performance on our bug analysis task compared to several other recent leading models.
To address RQ3 regarding the generalizability of BugLens, we also tested its performance with a range of prominent alternative models, encompassing both closed-source and open-source options. These include: OpenAI’s o1 (o1-2024-12-17), GPT-4.1 (gpt-4.1-2025-04-14), Google’s Gemini 2.5 Pro (gemini-2.5-pro-preview-03-25), Anthropic’s Claude 3.7 Sonnet (claude-3-7-sonnet-20250219), The open-source DeepSeek R1 (671B) model. This selection allows us to assess how BugLens performs across different model architectures, sizes, and providers.
5.2.1. Evaluation Dataset: Full Kernel Analysis with static analyzers
Our study utilizes the Android kernel (Linux version 4.14.150, Google Pixel 4XL), which served as the testbed in the original Suture study (zhang_statically_2021, ). For our initial analysis, potentially informing RQ1 regarding baseline performance and the challenges of automated bug detection in this complex environment, we applied prior static analysis tools, Suture and CodeQL-SOD, to the Linux kernel device drivers. The key findings from this analysis provide important context:
  • $\bullet$
    Suture: When applied to the Linux device drivers, Suture initially generated 251 potential bug reports. This raw output translates to a high False Positive (FP) rate of approximately 90%. Suture employed a subsequent semi-automated refinement process, reporting a reviewer-perceived FP rate of 51.23%. However, this figure relies on Suture’s broader definition of a bug, which classified all integer overflows as true positives. Furthermore, during our investigation, we identified 4 additional instances, initially dismissed as FPs by Suture authors’ verification, that were indeed real bugs (by either their standard or ours). This finding highlights potential inconsistencies in large-scale manual verification efforts.
  • $\bullet$
    CodeQL-SOD: We leverage CodeQL (codeql, ) as a complementary static analysis tool beyond Suture. Based on Backhouse et al.’s approach (backhouse_stack_2018, ), we implemented a simple inter-procedural taint tracking analysis (CodeQL-SOD) that traces data flows from ioctl entry points to copy_from_user sink functions to detect potential stack overflows. Running CodeQL-SOD on the Linux device drivers, it yields 11 potential bug reports. Our manual analysis confirmed 1 of these as a true positive (consistent with the known stack exhaustion bug reported in (backhouse_stack_2018, )). This corresponds to an FP rate of 90.9%.
5.2.2. Cost.
On average, the cost of running BugLens is about \$0.1 per case, under the latest version of OpenAI o3-mini. Each case takes about a few minutes to complete.
5.3. RQ1: Effectiveness
Table 2. Performance of BugLens on top of Suture and CodeQL-SOD.
Method TP TN FP FN Prec Rec F1
Suture 24 0 227 0 0.10 1.00 0.17
SutureRP 20 202 25 4 0.44 0.83 0.58
SutureBugLens 24 218 9 0 0.72 1.00 0.84
CodeQL-SOD 1 0 10 0 0.09 1.00 0.17
CodeQL-SODBugLens 1 9 1 0 0.50 1.00 0.67
5.3.1. Precision and Recall.
Table 2 shows the performance for the evaluated two static analyzers and our post-refinement method, BugLens. The results show BugLens significantly enhances precision of both Suture (0.10) and CodeQL-SOD (0.09). For CodeQL-SOD, the refinement (CodeQL-SODBugLens) increases precision substantially to 0.50 while not missing any real bugs detected before. For Suture, the precision is increased to 0.72 due to a drastic reduction in false positives (from 227 to 9). This refinement does not miss any exisiting bugs.
Moreover, noting that the semi-automated method in Suture, noted as SutureRP, actually shows a lower recall (0.83) than the BugLens refinement (1.0). This is because after examining the positive results of SutureBugLens, we found 4 cases of real bugs that were incorrectly classified as false positives during the manual inspection process in SutureRP. The results for SutureRP are estimated on the performance reported from its paper (zhang_statically_2021, ), as the semi-automated approach involves manual analysis steps that we did not replicate.
5.3.2. New Bugs.
As mentioned in the previous part, we find 4 more cases that are actually real bugs, which are previously classified as false positive by human inspection in Suture. Two bugs are from the sound subsystem, reported to the maintainers, while waiting for their feedback. One of them involves a data constraint that appears to sanitize a tainted value but can be bypassed due to subtle control-flow logic.
The other two bugs are from the i2c subsystem. They involve a condition where two tainted values must simultaneously satisfy specific constraints—a case that standard taint analyses typically miss due to their focus on single-source propagation.
We have reported all cases following responsible disclosure. Full technical details will be made available after the issues are resolved.
5.3.3. Analysis of FPs.
Despite the general effectiveness of BugLens, it shows 10 FPs. Upon careful examination of these cases, we attribute the inaccuracies to several distinct factors:
  • $\bullet$
    Static Analysis Fundamental Limitations (5 cases): False positives arising from inherent limitations in the underlying static analyzers that BugLens is not designed to address. These include imprecisions in taint tracking through complex data structures, incorrect indirect call resolution, etc. BugLens intentionally operates on the dataflow provided by the static analyzers rather than attempting to verify the accuracy of this information itself.
  • $\bullet$
    Environment and Language Understanding (4 cases): Imprecision resulting from the LLM’s incomplete grasp of C language semantics, hardware-level interactions, and kernel-specific programming patterns.
  • $\bullet$
    Internal Modeling Errors (1 case): Inaccuracy originating from a faulty prediction by an internal analysis component (LLM).
5.3.4. Analysis of FNs.
Despite the number of FN shown in the table 2 is 0, BugLens still produced several FNs but gets mitigated by majority voting. Specifically, we observed that the FNs were concentrated in the Constraint Assessor (ConA) component, and the SecIA component typically does not generate FNs due to the arbitrary control hypothesis (AC-Hypo), which will be discussed in §5.4.2 (we also provide a case study in §A).
We identified two primary reasons for these FNs:
  • $\bullet$
    Overlooked Bypass Conditions: The LLM sometimes failed to recognize complex conditions allowing mitigations or checks within the code path to be bypassed, as discussed in §3.3.5.
  • $\bullet$
    Misinterpreting Validation vs. Sanitization: LLMs occasionally misclassified sanitization as validations, thereby missing vulnerabilities, as detailed in §3.4.
These specific failure patterns observed within ConA were the direct motivation for designing the Structured Analysis Guidance (SAG). The SAG mechanism enhances the prompts used specifically within the ConA stage, providing targeted instructions aimed at guiding the LLM to avoid these identified pitfalls (e.g., explicitly probing for bypass logic, carefully differentiating data constraint types).
The positive impact of SAG in mitigating these FNs is empirically demonstrated in RQ2 (§5.4). As shown in Table 3, the Full Design configuration (using ConA with SAG prompts) consistently yields fewer FNs.
Nevertheless, SAG represents guidance, and its effectiveness is not absolute. The degree to which an LLM can accurately follow these complex instructions can vary, especially across different models. Our results (RQ2/RQ3, Table 3) indicate that while SAG provides significant benefits, some models may still struggle to fully leverage the guidance, potentially resulting in some residual FNs.
Takeaway. BugLens can effectively post-refine the results of exisiting static analyzers, hugely improving the precision, and can even find missed bugs.
Table 3. Bug Analysis Performance Comparison Across LLMs and Design Approaches (Total Cases=120, Real Bugs=22)
Model Full Design w/o SAG Baseline
FN FP F1 FN FP F1 FN FP F1
[…]
Gemini 2.5 Pro 12 3 0.57 14 4 0.47 6 24 0.52
Claude 3.7 Sonnet 13 2 0.54 17 2 0.34 1 51 0.44
[…]
Table 4. Performance of SecIA, with and without the Aribitary Control Hypothesis (AC-Hypo).
Model w/o AC-Hypo w/ AC-Hypo
FP FN Prec Rec FP FN Prec Rec
OpenAI o3-mini 13 5 0.57 0.77 38 0 0.37 1.0
OpenAI o1 8 4 0.69 0.82 39 0 0.36 1.0
OpenAI GPT-4.1 15 2 0.57 0.91 36 1 0.69 0.95
Gemini 2.5 Pro 60 3 0.24 0.86 73 0 0.23 1.0
Claude 3.7 Sonnet 20 2 0.50 0.91 31 0 0.42 1.0
DeepSeek R1 25 12 0.29 0.45 71 4 0.18 0.82
5.4. RQ2: Component Contribution
To address RQ2, we conduct an incremental analysis. This study evaluates the contribution of the Security Impact Assessor (SecIA), the subsequent Constraint Assessor (ConA), and the specialized Structured Analysis Guidance (SAG) design used within ConA, by comparing performance across progressively enhanced configurations of BugLens.
We assess the performance under the following configurations:
  • $\bullet$
    Baseline: This configuration employs the simple prompt design (§3.1.2) without the SecIA or ConA components. It establishes the baseline performance relying primarily on the LLM’s inherent capabilities with minimal guidance. The baseline design provides a starting point for comparison.
  • $\bullet$
    + SecIA: Adds the SecIA component to the Baseline. Purpose: Comparing this to the Baseline isolates the contribution of the SecIA stage. Detailed metrics for SecIA’s filtering rate and soundness are in Table 4.
  • $\bullet$
    + SecIA + ConA (w/o SAG): Adds the ConA component to the ”+ SecIA” configuration, utilizing a simpler prompt design for constraint checking (i.e., without SAG). Comparing this to “+ SecIA” isolates the contribution of adding the constraint assesses step itself.
  • $\bullet$
    Full Design (+ SecIA + ConA + SAG): This configuration enhances the ConA component from the previous step by incorporating the specialized SAG design. This represents the complete BugLens system. Comparing this to previous configurations emphasize the contribution of the SAG.
For this component analysis (RQ2) and the subsequent model versatility analysis (RQ3), we focus our evaluation on a dataset derived from Linux kernel analysis, specifically targeting the sound module. This module was selected because the original Suture study identified it as containing a high density of true positive vulnerabilities (22 out of 24 known bugs), providing a rich testbed for assessing bug detection capabilities. The dataset consists of 120 cases, with 22 known bugs (positives) and 98 non-bug cases (negatives).
We evaluate the performance of these components for diverse LLMs, including OpenAI’s o3-mini, o1, GPT-4.1, Gemini 2.5 Pro, Claude 3.7 Sonnet, and DeepSeek R1. The overall performance results for these configurations are summarized in Table 3, while Table 4 provides the detailed breakdown specifically for the SecIA component’s effectiveness and filtering metrics.
5.4.1. Baseline.
Our experimental results clearly demonstrate the significant contribution of our proposed multi-phase workflow and its components compared to a baseline approach. As shown in Table 3, despite some models like Claude 3.7 Sonnet (FN=1) and GPT-4.1 (FN=3) showed low False Negatives, potentially reflecting their raw analytical power, this came at the cost of high False Positives (FP=51 and FP=23, respectively), rendering this simple design ineffective for practical use. The F1 scores for the baseline were generally low across models. This direct prompting approach demonstrated worse performance when compared to other BugLens configurations.
5.4.2. Security Impact Assessor (SecIA)
As Table 4 shows, our Arbitrary Control Hypothesis (AC-Hypo) enhances recall across all models. Without AC-Hypo, the models exhibit noticeable False Negatives, ranging from 2 to 12 FN cases across tested models. After applying AC-Hypo, the FN rate decreases to zero for all models except DeepSeek R1 and GPT-4.1, which maintains a low FN rate (4 and 1, respectively), achieving a high recall of 0.82 and 0.95.
Meanwhile, SecIA demonstrates strong effectiveness as a security vulnerability filter. Taking OpenAI’s o3-mini as an example, among a total of 98 negative cases, SecIA successfully filters out 60 cases (TN). This indicates that SecIA not only has a high recall rate, but it is also effective, substantially improving analysis efficiency.
5.4.3. Constraint Assessor (ConA) without Structured Analysis Guidance (SAG)
While this multi-phase workflow (i.e., SecIA + ConA) significantly reduces the high volume of FPs seen in the Baseline; for instance, Claude 3.7 Sonnet’s FPs dropped from 51 to 2, and Gemini 1.5 Pro’s from 24 to 4. It also leads to a significant increase in False Negatives (FNs) for Gemini 2.5 Pro (FN=14), Claude 3.7 Sonnet and DeepSeek R1 (FN=17), and OpenAI o1 (FN=10) in the ‘w/o SAG’. This supports our hypothesis (§3.1.2) that providing constraints, while helpful for pruning obvious non-bugs, can encourage LLMs to become overly confident. Once patterns suggesting data validity are identified, the LLM may default to classifying the issue as “not a bug,” reflecting a potential statistical bias towards common safe patterns rather than performing nuanced reasoning about subtle flaws or bypass conditions.
5.4.4. Structured Analysis Guidance (SAG)
Comparing the Full Design (using SAG within ConA) to the w/o SAG configuration in Table 3 demonstrates SAG’s effectiveness. Introducing SAG leads to a substantial reduction in FNs across all tested models. Consequently, the overall F1 score sees a marked improvement with SAG (e.g., improving from 0.67 to 0.94 for o3-mini and 0.34 to 0.54 for Claude). This indicates that SAG successfully guides the LLM within ConA to overcome the previously observed overconfidence, achieving a better balance between FP reduction and FN mitigation.
Takeaway. The design components of BugLens enables effective LLM bug analysis by significantly reducing both FP and FN compared to baseline prompting.
5.5. RQ3: Model Versatility
The results shown in Table 3 affirm that BugLens is a general LLM-based technique applicable across different models, consistently improving upon baseline performance. However, the degree of success highlights variations in how different LLMs interact with complex instructions and structured reasoning processes.
As noted in RQ2, the baseline performance offers a glimpse into the models’ raw capabilities, somewhat correlating with general LLM benchmarks where Gemini 2.5 Pro and Claude 3.7 Sonnet are often considered leaders (vellum2025leaderboard, ). However, this raw capability did not directly translate to superior performance within our structured task without significant guidance.
When employing the ‘Full Design’ of BugLens, we observed distinct differences in instruction-following adherence. The OpenAI models, o1, GPT-4.1 (F1=0.81), and particularly our core model o3-mini (F1=0.94), demonstrated excellent alignment with the workflow’s intent.
Conversely, while the ‘Full Design’ significantly improved the F1 scores for Gemini 2.5 Pro (0.57) and Claude 3.7 Sonnet (0.54) compared to their baseline or ‘w/o sag’ results by drastically cutting down FPs, they still struggled with relatively high False Negatives (FN=12 and FN=13, respectively). This suggests that even with the SAG, these powerful models may face challenges in precisely balancing the various analytical steps or interpreting the nuanced instructions within our workflow, possibly still exhibiting a degree of the previously mentioned over-confidence (for “sanity check”) that SAG could not fully overcome in their case. DeepSeek R1 (F1=0.77) showed a strong, balanced improvement, landing between the OpenAI models and the Gemini/Claude in terms of final performance with the full design. This demonstrates the potential of the BugLens using in open-sourced models.
In summary, while our approach is broadly applicable, its optimal performance depends on the LLM’s ability to robustly follow complex, multi-step instructions, with models like OpenAI’s o3-mini currently showing the strongest capability in this specific structured bug analysis task.
Takeaway. BugLens shows broad applicability and yields promising results across diverse LLMs, including the open-source DeepSeek R1. OpenAI’s o3-mini currently gets the best result.
5.6. Case Study: Data Structure Traversal
Linux kernel code often uses pointers to traverse data structures. For example, the following code walks through a linked list list using a marco list_for_each_entry:
In this code, the loop goes through each element in the list and checks if any of them match the input id, which comes from the user. This kind of loop is hard for static analysis tools to understand unless they are specially designed to handle linked lists. These tools might wrongly report a warning, saying that user input affects when the loop ends. But in reality, the loop always stops at the end of the list, not because of the user input.
The user input only decides which element gets picked, not how long the loop runs. Our method, BugLens, avoids this false warning by using the LLM’s deeper understanding of how data structures like linked lists work.
5.6.1. A Bypassable Condition
There’s also a tricky part in this code. Before the list_for_each_entry loop even runs, there is a condition that checks if the id->numid is not zero. If it is not zero, the function will return the result of snd_ctl_find_numid(card, id->numid).
This means the checks inside the loop for id->index might not happen at all. So even the loop contains conditions seem to validate id->index, those checks is not always are only performed if id->numid is zero. Otherwise, the function skip the loop entierly.
This cases directly causes false negatives in BugLens, especially without the SAG mechanism. Limited by space, we describe this case in detail in the appendix §A.1.
6. Limitation & Discussion
Data Bias. The potential data bias introduced by Suture and CodeQL-SOD could threaten the internal validity. Despite the design of BugLens is completely decoupled from them, we only make tests on top of them might introduce bias into our experiments. BugLens may perform differently on other static analysis tools.
Towards More Sound Analysis. The soundness of current implementation could be improved through two directions: (1) Adding symbolic verification (bhatia_verified_2024, ; cai_automated_2025, ) to validate the LLM’s reasoning and refine outputs based on formal methods (2) Implementing a hybrid architecture where the LLM performs initial code slicing while symbolic execution handles constraint analysis, combining the LLM’s contextual understanding with provably sound formal reasoning.
Intergation with More Static Analysis Tools. The CodeQL-SOD is an interesting static analysis tool that is based on the lightweight static analysis framework, CodeQL. Considering that BugLens can effectively improve the precision of existing static analysis tools, we could implement more corsare-grained (with low precision) static analysis tools and integrate them into BugLens. Compared to heavy static analysis tools such as suture, an imprecise static analysis could be much easier to implement and maintain.
7. Related work
LLMs for bug detection. LLMs have shown promise in bug detection tasks, leveraging their ability to understand code semantics and context. Recent approaches like LLM4SA (wen_automatically_2024, ), LLift (llift, ), and LLMSAN (DBLP:conf/emnlp/WangZSX024, ) combine static analysis with LLMs to refine results or enable end-to-end bug detection. BugLens takes a different approach by leveraging LLMs to reason about complex path constraints and semantic conditions, addressing scenarios where traditional symbolic methods struggle.
LLMs for program analysis. LLMs have been widely applied to program analysis tasks, including static semantics analysis (DBLP:conf/nips/WangZSXX024, ; DBLP:journals/corr/abs-2501-18160, ), indirect call resolution (cheng_semantic-enhanced_2024, ), and various inference tasks in program verification and synthesis  (chow_pyty_2024, ; le-cong_can_2025, ; pei_can_2023, ; misu_towards_2024, ; cai_automated_2025, ; li_iris_2025, ). Similarly, BugLens leverages LLMs to reason about security impacts, utilizing their contextual understanding of code semantics to enhance the precision of program analysis results.
Reasoning for LLMs. Despite their success on many tasks, the ability of LLMs to reason about code semantics and behaviors remains an active area of research (ding2024semcoder, ; zelikman2022star, ; li2025codei, ; ni2024next, ). Recent studies have shown that LLMs are still far from performing reliable code reasoning, and their predictions are thus fragile and susceptible to superficial changes in input (pei2023exploiting, ; steenhoek_err_2025, ; hochlehnert2025sober, ). This fragility is often attributed to the learned models taking “shortcuts” based on superficial patterns in training data rather than robust, generalizable reasoning strategies (mccoy_embers_2024, ; yefet2020adversarial, ; gao2023discrete, ; yang2022natural, ; bielik2020adversarial, ). BugLens mitigates this problem with a similar spirit to existing works on boosting the LLMs’ reasoning by constraining their reasoning space with structural and symbolic procedures (li2025structured, ; chen2022program, ; li2023chain, ; chen2024steering, ; yao2023tree, ).
8. Conclusion
This paper introduces BugLens, an innovative post-refinement framework that integrates Large Language Models (LLMs) with static analysis. By employing Security Impact Assessor (SecIA), Constraint Assessor (ConA), and Structured Analysis Guidance (SAG) to guide LLMs through the reasoning process, BugLens significantly enhances the precision of initial static analysis findings without sacrificing scalability. Our evaluation demonstrates that BugLens dramatically reduces false positives in Linux kernel Analysis, minimizes manual inspection effort, and uncovers previously ignored vulnerabilities, highlighting the promise of guided LLMs in making automated bug detection more practical and effective.
Appendix A Case Study of BugLens
A.1. Data Structure Traversals & Bypass Conditions
In §5.6, we have shown a case of data structure traversal:
We now consider a caller of our example snd_ctl_find_id. The function snd_ctl_elem_write takes a control parameter, which is also tainted user input. It then calls snd_ctl_find_id to find the corresponding kctl object. The put function of kctl is then called with the tainted control parameter.
A.1.1. Step 2: Constraint Collection
Suppose the put function is a sink or critical operations that could lead to a vulnerability (and there are indeed many such cases, for different instances of kctl).
When Constraint Assessor (ConA) analyzes this code, its second step is to collect the data constraints of for the tainted input control (actually, the tainted input is control->id, and specific field depends on how the exactly put uses).
Natrually, it will collect the following constraints: snd_ctl_find_id(...) cannot return NULL. and the kctl object must be valid.
A.1.2. Step 3: Constraint Effect Summarization
After collecting the constraints, the next step is to summarize the effect of the constraints, so we need to go back to the snd_ctl_find_id function and analyze the constraints of control->id.
Here’s the most triky part, suppose we track the taint data control->id.index and by looking at the definition of snd_ctl_find_id, we can see that:
Naively, LLM may think these two conditions ensures the control->id.index is within a strict range, between kctl->id.index and kctl->id.index + kctl->count Considering that these kctl objects are predefined and maintained in the kernel, we might think our tainted input is used with an effective and restricted range.
A Closer Look: However, this is not the case. The key here is before the list_for_each_entry loop:
Before the loop, the function checks if id->numid is not zero, and if so, it calls snd_ctl_find_numid to find the corresponding kctl object, and returns it directly. This means that the list_for_each_entry loop, and inside checks can be bypassed if the id->numid is not zero.
The function snd_ctl_find_numid is defined as:
This function also traverses the card->controls list, but it uses kctl->id.numid as the key to find the corresponding kctl object. The numid is a unique identifier for each kctl object, and it is not directly related to the control->id.index field.
Finally, the current effect summarization for the snd_ctl_find_id function should be like (not shown completely):
  1. (1)
    Precondition: control->id.numid != 0,
    Postcondition: kctl != NULL && kctl->id.numid <= control->id.numid < kctl->id.numid+kctl->count
  2. (2)
    Precondition: control->id.numid == 0
    Postcondition: kctl != NULL && kctl->id.index <= control->id.index < kctl->id.index+kctl->count
The overlook of this bypass is the main reason of the false negatives for models.
Appendix B Outline of LLM Prompt
We provide the outline of the LLM prompt for each task in BugLens. For simplified, we remove all examples (that used in few-shot in-context learning).
The args are the arguments for each case running, and the callback is the callback function for each prompt for the agentic design PKA.
B.1. Prompt for Inferring Variable Names
I have a static analysis tool that tracks tainted user input in Linux kernel drivers.
Since the analysis uses LLVM IR, I need help identifying the corresponding source-level variables.
For each case:
1. Tainted values are either local variables loaded from globals/parameters, or direct function parameters
2. Tainted values are always numeric
3. Be aware of name redundancy, especially in global variables and their fields
4. Identify specific struct field names when applicable
5. For propagation, identify the first local variable receiving the taint
Analyze line by line:
The bug detector is: {}
source code:
Instructions:
line no (in source code): {}
args:
- get_bug_detector
- get_function_first_part
- get_insts_from_ctx
- get_source_line_set
B.2. Prompt for SecIA
I have a static analysis tool for Linux kernel drivers that produces many false positives when detecting:
1. Tainted Arithmetic Operations (integer overflows)
2. Tainted Loop Bound Conditions (infinite loops, unexpected iterations)
3. Tainted Pointer Dereferences (arbitrary memory access)
4. Buffer Overflow (out-of-bounds access)
5. Tainted length in copy_from_user (especially stack overflow)
Your Task:
1. Analyze the provided code snippet flagged by our tool
2. Assume:
- Attackers cannot control the kernel
- Tainted variables can be set to any value within their type range
- All existing checks can be bypassed
- Ignore any security checks in the code
3. Determine if the code represents:
- Potential Bug: If the tainted variable can cause infinite loops, very large loops, or memory bugs in the current context
- Normal Code: If the tainted variable usage doesnt lead to security issues within our scope
Provide a step-by-step explanation of your reasoning and classification.
tainted_varaible: {}
bug detector: {}
source code
args:
- get_tainted_value
- get_bug_detector
- get_function
B.2.1. Schema-constrained summarization
Summarize our discussion, and respond with a <bug_eval> tag indicating whether the tainted variable can lead to vulnerabilities:
currently we only consider (1) infinite/very-large loop, (2) out-of-bound access/buffer overflow/arbitrary memory access;
for other types of bugs, you can say "normal code" or "not_a_bug"
<bug_eval>
<tainted_var>tainted_var</tainted_var>
<vulns>
<vuln>
<type>out_of_bound_access</type>
<desc>Brief description of how the vulnerability occurs</desc>
</vuln>
</vulns>
</bug_eval>
For no vulnerability: <bug_eval>not_a_bug</bug_eval>
For potential vulnerability: <bug_eval>potential_bug</bug_eval>
If uncertain: <bug_eval>uncertain</bug_eval>
B.3. Prompt for ConA
This part we provide the prompt for the Constraint Assessor (ConA) task. These prompts already contain the design of SAG. Additionally, we also provide a list of functionalities of the agent PKA inside prompts, represented as {AGENT PROMPTS HERE}.
B.3.1. Step 1: Reachability Analysis
The following is the prompt for the step 1 of ConA task. (see §3.3.3 for details)
Identify the sink (the last line of the provided function context) and determine its preconditions.
Focus only on two types of preconditions:
1. Direct checks (dominate conditions) - Conditions that directly control if the sink executes
Example: if (flag) { sink(tainted_var); }‘
2. Early returns/bypasses (guard conditions) - Conditions that cause early returns or jumps that bypass the sink
Example: if (tainted_var > 100) return; sink(tainted_var);‘
Here, "tainted_var <= 100" is the precondition to reach the sink.
Ignore conditions that dont directly impact sink reachability.
{AGENT PROMPTS HERE}
sink variable
sink context (the full context of the function):
Summarize in this format:
‘‘‘xml
<sink_precondi>
<precondi>
<type> dominate_condition </type>
<condition> flag </condition>
<dominated_sink> if(flag) sink(tainted_var) </dominated_sink>
</precondi>
<precondi>
<type> guard_condition </type>
<condition> tainted_var <= 100 </condition>
<guard_bypass> if (tainted_var > 100) return/goto invalid_label; </guard_bypass>
</precondi>
</sink_precondi>
Note: Multiple preconditions combine with "AND" logic.
args:
- get_tainted_value
- get_function
callback:
- need_struct_def
- need_global_var_def
B.3.2. Step 2: Constraint Collection
The following is the step 2 of the prompt for the ConA task. (see §3.3.4 for details)
Help find range constraints for the tainted variable across the call chain.
tainted_variable: {}
current callchain: {}
context of the sink:
Focus on finding all possible value range constraints for the tainted variable, without analyzing their effectiveness.
Types of constraints to look for:
1. Validation: conditions that reject invalid ranges (e.g., if (tainted_var < 0) return -EINVAL;)
2. Sanitization: corrections applied to the value (e.g., tainted_var = min(tainted_var, 100);)
3. Type constraints: implicit limits from variable types (e.g., uint8_t limits to [0,255])
Important notes:
- Validations are transferable through operations (e.g., constraints on var = tainted_var + 1 apply to tainted_var)
- Sanitizations are not transferable
- Base your analysis only on the provided code, not prior knowledge
- Track the variable across different names in different contexts
{AGENT PROMPTS HERE}
Summarize your findings in this format:
<range_constraints>
<constraint>
<type>validation|sanitization|type_constraint</type>
<handler_func>function_name</handler_func>
<context>relevant_code_snippet</context>
</constraint>
</range_constraints>
args:
- get_tainted_value
- get_call_chain
- get_function_first_part
callback:
- need_func_def
- need_caller
- need_struct_def
- need_global_var_def
B.3.3. Step 3: Constraint Effect Summarization
The following is the prompt of the step 3 of the ConA task. (see §3.3.5 for details)
Act as a program verifier/symbolic execution engine to infer precondition and postcondition pairs for constraints in code. Use these simple rules:
- The **precondition** is the path condition required to reach the constraint.
- The **postcondition** is the effect on the tainted variable (often shown as its valid range).
Consider all branches and early return "bypass" cases. For instance, given this sample function:
‘‘‘c
void check_tainted_value(int config, int other_config, int x){
if (other_config == CHECK_SKIP)
return;
if (config == 1)
assert(x > 0 && x < 100);
else if (config == 2)
assert(x > 0);
You must extract the following [precondition, postcondition] pairs:
1. Bypass Case
Precondition: other_config == CHECK_SKIP
Postcondition: x in (-inf, +inf)
2. Config 1 Case
Precondition: other_config != CHECK_SKIP && config == 1
Postcondition: x in (0, 100)
3. Config 2 Case
Precondition: other_config != CHECK_SKIP && config == 2
Postcondition: x in (0, +inf)
4. Default Case
Precondition: other_config != CHECK_SKIP && config != 1 && config != 2
Postcondition: x in (-inf, +inf)
{AGENT PROMPTS HERE}
Finally, output your analysis as XML in the following format:
‘‘‘xml
<range_constraint>
<type>validation</type>
<handler_func>check_tainted_value</handler_func>
<condition_pairs>
<pair>
<precondi>other_config == CHECK_SKIP</precondi>
<postcondi>x in (-inf, +inf)</postcondi>
<context> if (other_config == CHECK_SKIP) return; </context>
</pair>
<pair>
<precondi>other_config != CHECK_SKIP && config != 1 && config != 2</precondi>
<postcondi>x in (-inf, +inf)</postcondi>
<context> induced by the other branches </context>
</pair>
</condition_pairs>
</range_constraint>
callback:
- need_func_def
- need_struct_def
- need_global_var_def
- need_caller
B.3.4. Step 4: Final Vulnerability Evaluation
The following is the prompt of the step 4 of the ConA task. (see §3.3.6 for details)
Evaluate if this bug is eliminated, not exploitable, or still vulnerable.
Extract concrete values for "size of" or "length of" in constraints and bug conditions.
Analyze whether the bug condition can be satisfied considering all constraints:
- Disregard constraints with postconditions that dont limit the tainted variable
- Disregard constraints for unrelated sinks
- For kernel-controlled conditions, determine reachability based on your knowledge
- If a precondition contains user-controlled variables, assume users can bypass it
A bug is "eliminated" only if:
- The postcondition restricts the tainted variable to a safe range that makes the bug impossible
- The precondition is always satisfied when the sinks precondition is true
If the postcondition isnt strong enough, assess exploitability assuming attackers can set any value.
{AGENT PROMPTS HERE}
callback:
- need_func_def
- need_struct_def
- need_global_var_def
- need_caller
B.3.5. Schema-constrained summarization
The following is the prompt for summarization of the ConA task, which is used to extract the final result of the vulnerability evaluation within a <final_res>.
Now Lets summarize our discussion, and respond in a <final_res> tag with the following format:
"still_a_bug", "eliminated", "likely_safe", "likely_unsafe", "not_exploitable" or "uncertain" within a <final_res> tag, e.g., <final_res>still_a_bug</final_res>
B.4. Prompt for PKA
The following is the exact prompt that shown as AGENT PROMPTS HERE in the previous parts.
1========================================
2First of all, you dont need to complete the task in your initial response. You can always ask for more information.
3
4When you need additional details, use the following format. In this case, dont reach a conclusion immediately - instead, request the information you need to perform a thorough analysis once you receive my response.
5--- request 1: ask for the function definition ---
6You could ask me for the definition of the function. in this case, you could respond with the following:
7<requests>
8 <request>
9 <name>need_func_def</name>
10 <args>
11 <arg>func_1</arg>
12 <arg>func_2</arg>
13 </args>
14 </request>
15</requests>
16--- request 2: ask for the struct definition ---
17You could ask me for the definition of the structure. in this case, you could respond with the following:
18<requests>
19 <request>
20 <name>need_struct_def</name>
21 <args>
22 <arg>struct_name_1</arg>
23 <arg>struct_name_2</arg>
24 </args>
25 </request>
26</requests>
27--- request 3: ask for the caller of the current function ---
28You could ask me for the caller for the current function. (Note: you can only request one caller at a time)
29 in this case, you could respond with the following:
30 <requests>
31 <request>
32 <name>need_caller</name>
33 <args>
34 <arg>current_function_name</arg>
35 </args>
36 </request>
37</requests>
38--- request 4: ask for the definition of global variables ---
39You could ask me for the definition of global variables. In this case, you could respond with the following:
40<requests>
41 <request>
42 <name>need_global_var_def</name>
43 <args>
44 <arg>global_var_1</arg>
45 <arg>global_var_2</arg>
46 </args>
47 </request>
48</requests>
49========================================