### v3

# The Hitchhiker’s Guide to Program Analysis, Part II: Deep Thoughts by LLMs

Abstract.
Static analysis plays a crucial role in software vulnerability detection, yet faces a persistent precision-scalability trade-off. In large codebases like the Linux kernel, traditional static analysis tools often generate excessive false positives due to simplified vulnerability modeling and over-approximation of path and data constraints. While Large Language Models (LLMs) demonstrate promising code understanding capabilities, their direct application to program analysis remains unreliable due to inherent reasoning limitations.
We introduce BugLens, a post-refinement framework that significantly enhances static analysis precision for bug detection. BugLens guides LLMs through structured reasoning steps to assess security impact and validate constraints from the source code. When evaluated on Linux kernel’s taint-style bugs detected by static analysis tools, BugLens improves precision approximately 7-fold (from 0.10 to 0.72), substantially reducing false positives while uncovering four previously unreported vulnerabilities. Our results demonstrate that a well-structured, fully-automated LLM-based workflow can effectively complement and enhance traditional static analysis techniques.
conference: ; May 2025; Earth
1. Introduction
Static analysis has long served as a cornerstone technique for identifying software vulnerabilities. By analyzing the code without dynamic execution, these techniques aim to detect various security weaknesses, such as buffer overflows and information leaks. However, static analysis tools often struggle to balance the trade-off between precision and scalability (gosain_static_2015, ; DBLP:journals/csur/ParkLR22, ).
More precise analysis, for example, symbolic execution (horvath_scaling_2024, ), can be computationally expensive and often infeasible for large codebases such as the Linux kernel (ubitect, ). Conversely, more scalable techniques sacrifice the precision for scalability, leading to a high number of false positives. For example, Suture (zhang_statically_2021, ), an advanced taint bug detection in the Android kernel shows a 90% raw false positive rate, which requires manual inspection of the results.
Specifically, its imprecision stems from two main issues:
  • $\bullet$
    Simplified Vulnerability Modeling. Static analyzers may often rely on heuristic simplified rules for vulnerability detection. For example, a static analyzer may flag every arithmetic operation as potentially overflowing. While this simplication might ensure no genuine vulnerabilities are missed, it inflates the number of false positives.
  • $\bullet$
    Over-Approximation of Path and Data Constraints. In order to avoid exponential path exploration, static analyzers often make coarse assumptions about whether a path is feasible or how data flows through the program. This over-approximation ensures analysis completes in a reasonable time, but it also flags numerous infeasible paths as potentially vulnerable, resulting in excessive false positives.
Recent advances in Large Language Models (LLMs) offer a promising avenue for overcoming these issues. Trained on vast amounts of code and natural language, LLMs exhibit remarkable capabilities in understanding code semantics, API usage patterns, and common vulnerability types (khare_understanding_2024, ; stoica_if_2024, ; fang_large_2024, ). By leveraging these broader insights, LLM-based approach might able to: (1) enhance vulnerability modeling by providing a more nuanced understanding of code semantics, and (2) refine path and data constraints by a selective analysis of semantically plausible paths and data flows.
However, LLMs are not a silver bullet for program analysis. Despite their semantic understanding capabilities, LLMs are not inherently equipped for the rigorous demands of program analysis (qian_lamd_2025, ; he_code_2024, ). Their reasoning proves brittle, particularly when confronted with the complex program dependencies crucial for security analysis (kambhampati_can_2024, ; chen2025reasoning, ; chapman_interleaving_2025, ). Indeed, our initial experiments confirm that naively applying LLMs to program analysis, for instance, by simply asking “Is this warning a true positive?”, yields highly unreliable results, frequently misclassifying vulnerabilities or failing to identify critical flaws. This is often because LLMs tend to fixate on surface-level code features, missing the critical dependencies that dictate program behavior and security properties, especially within intricate control and data flows.
Our work addresses this challenge by introducing a structured guidance framework that directs LLM reasoning according to static program analysis paradigm. This approach rests on the premise that static analysis, while theoretically sound, is limited by practical tradeoffs in precision and scalability. By situating LLM reasoning within this established methodology, we “compel” the model to analyze more rigorously than it would by default, thus mitigating the inherent limitations of LLMs in code reasoning.
In this work, we introduce BugLens, an innovative framework that post-refines the results of static analysis using LLMs. Rather than blindly applying LLMs to analyze program, BugLens is carefully orchestrated to teach LLMs key concepts of program analysis and guide them toward reasonable analytical procedures. By analyzing the output of static analysis, BugLens  complements the limitations of existing tools, especially in terms of precision, and yields more accurate and actionable vulnerability detection for practical codebases. We demonstrate that this combined approach significantly improves the precision of taint-bug detection in the Linux kernel, reducing the need for manual inspection of false positives, and even uncovering previously ignored vulnerabilities.
We summarize our contributions as follows:
  • $\bullet$
    Post-Refinement Framework. We introduce BugLens, a post-refinement framework that supplements traditional static analysis to boost precision of the results, overcoming various practical weaknesses identified from real-world static analysis tools.
  • $\bullet$
    Structured Analysis Guidance (SAG). We design a structured workflow that directs LLMs to follow the principles and processes of established static analysis paradigms. This guided approach achieves better vulnerability detection compared to naive LLM prompting strategies.
  • $\bullet$
    Empirical Results. Evaluated on the Linux kernel, our solution improved the precision of a real-world static analysis tool from 0.1 to 0.72. Interestingly, prior manual analysis incorrectly filtered four warnings that were retained by BugLens.
  • $\bullet$
    Open Source. We open-source our code and data of BugLens on GitHub [^1] https://github.com/seclab-ucr/BugLens.
2. Background & Motivation
2.1. Taint Bugs
Taint-style bugs involve insecure data propagation, where unsanitized data (taint source) reaches a sensitive program location (taint sink), causing flaws like out-of-bound access. For instance, an unchecked input integer (source) might become an array index (sink). Proper sanitization (e.g., range checks) is typically missing or insufficient. Such bugs can lead to severe vulnerabilities like buffer overflows or denial-of-service.
In the kernel, these bugs often occur when untrusted user input (e.g., syscall arguments) propagates to sensitive kernel operations (e.g., arithmetic calculations) without adequate sanitization (machiry_dr_2017, ; zhang_statically_2021, ). Detecting these bugs is challenging due to:
  • $\bullet$
    Diverse and General Sinks. Unlike user-space applications with often well-defined sink APIs (e.g., exec()), the kernel’s high privilege means almost any operation (even a single arithmetic calculation) could be a sink if affected by unsanitized user data. This broad scope increases false positives. Additionally, low-level coding optimizations common in the kernel (e.g., intentional integer overflows) can complicate accurate detection.
  • $\bullet$
    Scattered and Intricate Sanitization. Sanitization checks can be distant from sinks, often crossing function boundaries within the large and complex kernel codebase. These checks are frequently intertwined with domain-specific kernel logic (e.g., privilege systems, configurations), posing significant hurdles for taint analysis tools.
2.2. Static Analysis for Taint Bugs
Suture (zhang_statically_2021, ) is the state-of-the-art work from academia aiming at kernel taint-style bug detection, which extends the previous work, Dr. Checker (machiry_dr_2017, ), by adding cross-entry taint tracking capability and improving analysis precision. CodeQL (codeql2025, ) is one of the most powerful and widely used industry code analysis engines, capable of taint analysis (taint_tracking_CodeQL, ) and query-based bug detection based on customizable rules. CodeQL has a large community that develops and maintains many different bug detection rules, including those for kernel taint-style bugs (e.g., (backhouse_stack_2018, )). Despite their effectiveness, all these tools exhibit high false alarm rate (e.g., ~90%), underscoring the inherent difficulties in accurately identifying kernel taint-style vulnerabilities. We elaborate on these difficulties and challenges in §3.
3. Challenges and Design Rationale
3.1. Challenges in Static Analysis
3.1.1. Challenge 1 (C1): Simplified Vulnerability Modeling
The first source of imprecision is the reliance on simplified vulnerability detection modeling. To maintain scalability and avoid missing potential bugs, static analyzers often employ overly simplified detection rules.
For instance, a static analyzer typically flags specific code patterns as potential bugs. A Tainted Arithmetic Detector (TAD) would identify any arithmetic operation involving tainted variables (e.g., var += size, where var is tainted) as a potential integer overflow vulnerability. However, in Linux kernel practice, these patterns frequently do not translate to exploitable vulnerabilities for two key reasons:
  • $\bullet$
    The Behavior Itself is Benign. The kernel’s reliance on low-level C idioms, pointer manipulation, and intentional ‘unsafe’ design patterns often confounds traditional analyzers, leading to inaccuracies, e.g., our evaluation shows that 61.2% of the flagged ‘bugs’ were actually benign (§6.4.2). For example, in Linux drivers, there are cases where integer overflow is expected and benign, such as with jiffies:
    For these cases, even when integer overflow is possible due to the tainted variable timeout, it is not a security vulnerability, as time_after is defined as #define time_after(a, b) (b - a) > 0). Such false positives commonly arise from intentional design patterns in the Linux kernel, including struct hacks (deliberate out-of-bounds access), type casting, unions, and data structure operations (e.g., ptr->next, container_of).
  • $\bullet$
    The Tainted Data is Properly Validated before reaching the sink. For example, the Taint in the Loop Bound (TLB) pattern is common in the Linux kernel, but many include proper validation. Consider this array traversal: [Uncaptioned image]
    Without the assert check, this loop could lead to an out-of-bounds access. However, the assert ensures that the loop index never exceeds the array bounds, making it safe. Static analyzers that does not account for such checks (e.g., path-insensitive ones) will misclassify this as a vulnerability, leading to false positives.
3.1.2. Challenge 2: Over-Approximation of Constraints
As mentioned above, the tainted data is often checked before the sink, and with a safe range. However, static analysis tools often over-approximate the path and data constraints to ensure scalability, as the kernel’s immense scale and continuous evolution make precise static analysis impractical, and the logic of the sanitization could often be distant or complex. Alternatively, a common approach in static analysis (infer_2015, ) is to directly prune paths with checks, regardless of their actual effects, which leads to false negatives. The existence of checks is not always effective in preventing the bug from happening, e.g., we discovered some unreported bugs due to a plausible but ineffective check (see §6.3). Consider the function clamp_num() shown in the following, which is designed to sanitize the tainted variable num:
The function clamp_num only sanitizes the tainted variable num effectively when the config is set. Therefore, a precise static analysis needs to perform (i) a range analysis on the tainted variable num, and (ii) a path-sensitive analysis on clamp_num() to determine path conditions.
3.2. The Opportunity in LLMs
Recent advances in Large Language Models (LLMs) offer a promising avenue to address these challenges, particularly in post-refining the results generated by scalable static analyzers. Trained on vast amounts of code and natural language, LLMs exhibit remarkable capabilities relevant to overcoming the limitations described above:
  • $\bullet$
    Enhancing Vulnerability Modeling (Addressing C1): LLMs possess a nuanced understanding of code semantics, common programming patterns, and API usage (li2025codei, ; ni2024next, ; liu_generating_2025, ; khare_understanding_2024, ). They can potentially recognize complex or subtle vulnerability indicators that are difficult to capture with predefined simplified heuristics. By analyzing code context (including comments and variable names), LLMs might differentiate between genuinely unsafe patterns and benign ones.
  • $\bullet$
    Refining Path and Data Constraints (Addressing C2): While not performing formal reasoning, LLMs can leverage their semantic understanding to assess the feasibility of paths and data flows flagged by static analysis. They might leverage contextual information to determine if the flagged bug is likely infeasible in practice, thereby helping to prune false positives arising from coarse path and data constraint analysis.
Recent works(wen_automatically_2024, ; li_iris_2025, ) also show that LLMs can be used to refine static analysis results (C1 & C2) in a straightforward manner, i.e., to directly ask LLM with “Is this case from static analysis report vulnerable?” followed by the code snippet.
3.2.1. Challenge 3: Reasoning Hurdles for LLMs.
Despite their potential, LLMs are not inherently suited for rigorous program analysis tasks out-of-the-box. Their reasoning can be fragile when dealing with complex program dependencies, and they lack systematic constraint-solving mechanisms. Our experiments (detailed in §6.4) demonstrate that such a simple prompting strategy leads to high false negatives — cases where the LLM incorrectly classifies actual vulnerabilities as safe code, especially against the presence of complex control flow and data constraints.
We observe that the false negatives stem from the fundamental limitation of existing LLMs when tasked with sophisticated, structured, and symbolic reasoning. Specifically, with simple prompting, LLMs often rely on surface-level spurious features (mccoy_embers_2024, ) to reason about the vulnerability. For example, the presence of a sanity check often correlates statistically with safe code (many practical static analyzers use such heuristics to prune paths). We observe the model classifies code containing such a check as safe without performing the deeper reasoning required to determine if the check is actually effective under the current execution paths. Real-world vulnerabilities often exploit exactly these scenarios: checks that are bypassable, incomplete, or rendered ineffective by intricate control and data flows. The model’s tendency to rely on spurious patterns without performing the actual reasoning has been shown ineffective in tracking critical dependencies and results in unreliable predictions in many different domains (prabhakar_deciphering_2024, ).
3.3. Design Rationale
To address challenges C1 through C3, we introduce BugLens, a fully automated, multi-stage LLM-based framework designed specifically as a post-refinement layer for static analyzers. The key design philosophy of BugLens is to scale the test-time compute of LLMs by allocating more tokens for precise reasoning of the intended program behaviors based on the model’s understanding of program semantics (snell2024scaling, ). Specifically, BugLens consists of the following specific key components, each targeting a specific challenge:
  • $\bullet$
    Security Impact Assessor (SecIA): Addressing Simplified Vulnerability Models (C1). Instead of relying solely on predefined patterns, BugLens leverages LLMs to analyze the potential security impact if tainted data identified by static analysis were completely controlled by an attacker. It then evaluates whether the tainted value could potentially lead to security vulnerabilities (e.g., memory corruption, denial of service (DoS)), as detailed in §4.1. By focusing on semantic consequences rather than simplified vulnerability modeling, SecIA provides a more precise assessment of security impact.
  • $\bullet$
    Constraint Assessor (ConA): Addressing Over-Approximation of Path & Data Constraints (C2). ConA uses LLMs to analyze whether data constraints in the code are sufficient to prevent potential vulnerabilities from being triggered. By tracing how tainted data is processed and constrained, ConA performs heuristic reasoning (e.g., to find sanitizations with names check_x) to evaluate the effectiveness of these safeguards. This approach is designed to handle the complexity and scalability challenges of systems like the Linux kernel, where traditional formal methods may lack precision or scalability.
  • $\bullet$
    Structured Analysis Guidance (SAG): Addressing Reasoning Hurdles for LLMs (C3). As described in §3.2.1, simply prompting LLMs for analysis often leads to spurious reasoning. To this end, SAG grounds LLMs’ reasoning in ConA with the scaffold that describes the typical program analysis workflow. At a high level, SAG employs specific prompts to describe key analysis principles and include in-context examples to demonstrate how to systematically dissect code, trace dependencies, and evaluate conditions, especially in complex scenarios. This guidance constrains the LLM towards a more rigorous and reliable analysis process. §4.3 elaborates on the design of SAG.
Figure 1. Overview of BugLens, showing (A) Security Impact Assessor (SecIA) first assesses the security impact of the potential bugs identified by static analysis, and then (B) Constraint Assessor (ConA) assesses the data constraints and evaluates if the bug is feasible. ConA is guided by (C) Structured Analysis Guidance (SAG) to reason on the code more effectively, and can interact with the (D) Project Knowledge Agent (PKA) to get information about the codebase on-demand.
[…]
Overview of our workflow, showing (1) static analysis and (2) LLM-based filtering
4. Design
BugLens first takes the static analysis report as input, which identifies the context of the potential bugs and the tainted data flow, and then performs the following steps to evaluate the potential bugs, as shown in Figure 1:
  • $\bullet$
    Security Impact Assessor (SecIA): This component evaluates the security impact of the potential bugs identified by static analysis. It identifies the Critical Operations (COp) that are influenced by the tainted data and classifies them as either Normal Code or Requires Constraint Analysis.
  • $\bullet$
    Constraint Assessor (ConA): This component performs a multi-step analysis to evaluate the feasibility of the potential bugs. It collects the path conditions and data constraints, summarizes them, and evaluates whether they are effective in preventing the vulnerability.
  • $\bullet$
    Structured Analysis Guidance (SAG): This component provides the scaffold to constrain the LLM through pre-defined code reasoning procedures to help it understand the code and analyze the constraints effectively.
  • $\bullet$
    Project Knowledge Agent (PKA): This component allows the LLM to access the codebase on-demand, enabling it to retrieve global codebase information.
Additionally, BugLens also adopts the common prompting techniques to scale the test-time compute, such as (1) Majority-vote querying, where we query the model multiple times and take the most common answer; (2) Chain-of-Thought (CoT) prompting (wei_chain–thought_2023, ); and (3) Schema-constrained summarization, where a follow-up prompt requests the model’s own output in a predefined XML format, making LLM’s response easy to parse.
4.1. Security Impact Assessor (SecIA)
The core insight behind SecIA is based on the fundamental evaluative question: What are the consequences if tainted data assumes arbitrary values?
4.1.1. Core Assumption and Rationale
SecIA operates on a fundamental assumption regarding attacker capability at the point of initial impact assessment: Aribitary Control Hypothesis (AC-Hypo). For a given program location $K$ where static analysis reported an operation $Op(v)$ involving tainted data $v$ (the sink), (1) we hypothesize that an attacker can control $v$ to take any value $v_{atk}$ (within the constraints of its data type), and (2) we provisionally ignore any effects of checks or path conditions (even explicit checks) encountered on analysis. In other words, we assume that the attacker can take any value to anywhere (successors of the sink node in the control flow graph).
This hypothesis enables SecIA to streamline analysis. SecIA focuses solely on potential security impact. This permits early filtering of findings based purely on the consequence, (safely) reducing subsequent analysis load. Critically, this approach defers the complex analysis of actual program constraints—including path feasibility, value ranges, and importantly, whether those protective checks or sanitizers really work. This deferral mitigates false negatives (FNs) by preventing premature dismissal of vulnerabilities due to reliance on potentially bypassable checks or inaccurate LLM constraint reasoning about their effectiveness. The effectiveness is shown in §6.4.2.
4.1.2. Security Impact Analysis
SecIA performs a forward influence analysis to identify the influenced critical operations (COp) that are potentially affected by the tainted data $v$ from the sink location $K$. It then filters out the benign operations that are not security-sensitive based on whether this operation could (potentially) result in memory bugs (e.g., out-of-bound access, arbitrary memory access) or DoS. For instance, the jiffies case in §3.1.1 will be attributed as “not a bug”.
Figure 2. The workflow of Constraint Assessor (ConA) with an example, where the x is tainted and op(x) is the critical operation (COP). The sanitization function check_x(..., x) is affected by the config (noted as c) and other_config (noted as oc).
[…]
This figure illustrates the multi-stage workflow of ConA.
4.2. Constraint Assessor (ConA)
The Constraint Assessor (ConA) aims to identify whether the bug of COp can be triggered by analyzing the data constraints. As shown in Figure 2, the Constraint Assessor involves a four-step workflow:
  • $\bullet$
    Step 1: Critical Operation Reachability Analysis. The process begins by determining the conditions under which program execution can reach the specific Critical Operation (COp) location. This establishes the base requirements for the vulnerability to be possible.
  • $\bullet$
    Step 2: Constraint Collection. Next, the analysis traces the tainted data flow path(s) backward from the COp towards the data’s source. Along this path, it identifies code segments—such as conditional statements, assertions, or calls to validation functions—that appear intended to act as constraints on the tainted data’s value or range before it is used at the COp.
  • $\bullet$
    Step 3: Constraint Effect Analysis. Each potential constraint identified in Step 2 is then analyzed in detail. This step aims to understand the constraint’s specific effect: Under what conditions (precondition) does the constraint need to satisfy, and what impact does it have on the tainted variable’s possible numerical range (postcondition)?
  • $\bullet$
    Step 4: Final Vulnerability Evaluation. Finally, it performs an evaluation to determine if the potential vulnerability can still be triggered. If yes, the case is classified as a “Potential Vulnerability.” Otherwise, if the LLM determines that the constraints effectively prevent the vulnerability from being triggered, the case is classified as “Eliminated.”
This workflow leverages LLMs to interpret code (mimicking a formal reasoning process), identify relevant patterns and constraints, and perform the reasoning required for each step.
We employ this LLM-based approach acknowledging the trade-offs of soundness and precision. ConA aims to overcome the potential precision limitations of conservative abstractions and achieve broader applicability in the complex Linux kernel. In §6.3 and §6.5, we thoroughly examine how our design in relaxing the strict over-approximation introduces only a reasonably small number of false negatives. In the following sections (§4.2.1 through §4.2.4), we elaborate on the specific design, rationale, heuristic considerations, and soundness implications inherent in each stage of this analysis process.
4.2.1. Step 1: Critical Operation Reachability Analysis
The analysis begins by determining the conditions required for program execution to reach the specific Critical Operation (COp) location previously identified by SecIA as potentially vulnerability. These reachability conditions form the base constraints that must be satisfied for the vulnerability to be triggerable via this COp.
Condition Representation: When analyzing path conditions, we allow the LLM to generate natural language summaries that capture the semantic meaning of complex reachability conditions. For instance, where traditional analysis might struggle, an LLM might identify a condition like: “The operation is only executed if init_subsystem() returned zero (success), AND the device state in dev->status equals STATUS_READY.”
The LLM is later asked (in Step 4) to evaluate these semantic constraints to assess the potential impact. The natural language representation allows the LLM to leverage its understanding of the code’s intent and context but introduces a higher degree of uncertainty in the final evaluation, which we must acknowledge.
4.2.2. Step 2: Backward Constraint Collection
Once the Critical Operation (COp) and its local reachability conditions are identified, the next step is to gather potential constraints imposed on the tainted data before it reaches the COp. This involves tracing the data flow path(s) for the tainted variable backward from the COp toward its source(s).
Example: In Figure 2, tracing back from critical_op(x) in caller(), the LLM identifies the call check_x(...) to be a potential constraint on x. It would query the PKA for the function definition and analyze its effect, as detailed in Step 3.
4.2.3. Step 3: Constraint Effect Analysis
After collecting potential constraining code segments (like the function check_x()) in Step 2, this step analyzes the effect of these segments on the tainted variable. The goal is to understand how different execution paths within these segments modify the possible range of the tainted variable and under which conditions (preconditions) those paths are taken.
This step prompts the LLM to first identify all major execution paths through the provided code segment (e.g., check_x()), and then considers the preconditions and postconditions (data constraints of the tainted data) of each path.
Example: In Figure 2, we ask the LLM to analyze the function check_x() to summarize its effects on the tainted variable x. Particularly, the path $\pi_{1}$ and $\pi_{4}$ are bypass paths, and check_x() will not effectively limit the range of x with these preconditions.
4.2.4. Step 4: Final Vulnerability Evaluation
This final step synthesizes the precondition and postcondition from the previous analyses to determine if the identified constraints effectively neutralize the potential vulnerability associated with the Critical Operation (COp). And then it classifies the reported vulnerability as either “Eliminated” (constraints effectively prevent the vulnerability happening) or “Potential Vulnerability” (no effective constraints).
[…]
4.3. Structured Analysis Guidance (SAG)
To support the fine-grained code reasoning steps in ConA, we employ Structured Analysis Guidance (SAG) to scale the test-time compute of LLMs with structured reasoning templates and few-shot examples. At a high level, SAG grounds the reasoning procedures of LLMs with typical program analysis steps by eliciting more reasoning tokens during inference. Specifically, SAG assists Constraint Assessor (ConA) with the following two types of analysis: (i) Guided Stepwise Vulnerability Analysis with step-by-step instructions that decompose the analysis of precondition and postcondition (as described in §4.2), and (ii) Guided Path Condition and Data Constraint Analysis to demonstrate how to analyze challenging data constraints and path conditions from code.
4.3.1. Guided Path Condition Analysis
In the analysis of the path condition in Step 1 and Step 3 of ConA4.2.1 and §4.2.3), we guide the LLM using prompts designed to extract path conditions from the source code surrounding the operation of interest. For example, consider the path leading to the op(x) call within the caller function:
SAG asks the LLM to identify and categorize:
  • $\bullet$
    Bypass Conditions: Identify conditional statements where taking a specific branch avoids the operation. The LLM is instructed to extract the condition and negate it to find the requirement for not bypassing the operation.
    Example: The condition x > 100 leads to invalid, bypassing critical_op(x). The negated condition required to proceed towards the OP is x $\leq$ 100.
  • $\bullet$
    Direct Conditions: Identify conditional statements where taking a specific branch is necessary to reach the operation along the current path. The LLM extracts the condition directly.
    Example: Reaching sink(x) requires entering the if block, so the condition is other_config $\neq$ SKIP.
[…]
4.3.2. Guided Data Constraint Analysis
In the analysis of data constraints in Step 2 and Step 3 of ConA4.2.2 and §4.2.3), We consider the following data constraints:
  • $\bullet$
    Type constraints. The variable’s static type already restricts its range (e.g., uint8 is always in the range of $[0,255]$).
  • $\bullet$
    Validation (transferable to source). The program tests the value and aborts or reports an error if the test fails, without modifying the value. Because the check refers to the current value, the knowledge gained from this check (e.g., “the value must be $\geq 0$ on the success branch”) also applies to all source variables that influenced this value in the data flow.
  • $\bullet$
    Sanitization (not transferable to source). The program writes a new, corrected value back to the variable (e.g., clamping it to a range). This operation severs the connection to the original value, so any property we learn afterwards applies only to the sanitized copy, not to the original source variables in the data flow.
The key difference is that validation knowledge travels backward along the data-flow graph, while sanitization overwrites the flow and stops the transfer.
[…]
Formal program analysis naturally distinguishes between validation and sanitization; however, LLMs often confuse these concepts (e.g., missing transferred validation, treating sanitization as validation). SAG emphasizes these differences with few-shot examples, and therefore improves the LLMs’ reasoning on data constraints.
4.4. Project Knowledge Agent (PKA)
BugLens leverages Project Knowledge Agent (PKA), a simple LLM code agent equipped with custom tools to automatically navigate the codebase and retrieve the related code context during the analysis. It starts the analysis with the context of the sink and explicitly requests more context (e.g., function definitions get_func_def, global variable definition get_var_def, struct layouts) if needed. PKA then retrieves the relevant code snippets iteratively until the model self-decides that it has sufficient context to complete the analysis.
We provide a set of request types, enabling the LLM to gather broader codebase information at any point. This design is both flexible and extensible: new request types can be supported by adding corresponding backend callbacks. PKA searches the relevant code snippets based on CodeQuery (kopathy_ruben2020codequery_2025, ). We implement PKA in 500 LOC of Python code.
5. Implementation
BugLens is implemented with approximately 7k tokens (30k characters) in prompts (detailed in §B) and 2k lines of Python code that manages API requests and codebase querying functionality.
5.1. Inferring Variable Names
Static analysis tools typically operate on compiler-generated IR code (like LLVM IR), which differs significantly from source code representation. Since BugLens analyzes source code directly, we need to map variables between these representations.
The IR code provides line numbers that correspond to locations in the source code, along with data flow information. However, compiler optimizations often create, eliminate, or transform variables, making this mapping non-trivial. For example, LLVM might convert a simple ternary expression into a complex bitwise operation.
We leverage LLMs to perform this variable name inference by providing them with: (1) the line number from source code and (2) the data flow from IR code. The LLM then identifies the corresponding source-level variables, effectively bridging the gap between IR-level analysis and source code understanding.
6. Evaluation
6.1. Research Questions
Our evaluation aims to address the following research questions.
  • $\bullet$
    RQ1: (Effectiveness) How effective is BugLens in identifying vulnerabilities?
  • $\bullet$
    RQ2: (Component Contribution) How does the the prompt design affect the performance of BugLens? especially for the SecIA and SAG component?
  • $\bullet$
    RQ3: (Model Versatility) How does the performance of BugLens vary across different LLMs?
6.2. Experimental Setup
We primarily evaluate BugLens using OpenAI’s o3-mini model, specifically o3-mini-2025-01-31. This model was chosen as our primary focus because, as demonstrated in our evaluation for RQ3 (Section 6.5), it achieved the best overall performance on our bug analysis task compared to several other recent leading models.
To address RQ3 regarding the generalizability of BugLens, we also tested its performance with a range of prominent alternative models, encompassing both closed-source and open-source options. These include: OpenAI’s o1 (o1-2024-12-17), GPT-4.1 (gpt-4.1-2025-04-14), Google’s Gemini 2.5 Pro (gemini-2.5-pro-preview-03-25), Anthropic’s Claude 3.7 Sonnet (claude-3-7-sonnet-20250219), The open-source DeepSeek R1 (671B) model. This selection allows us to assess how BugLens performs across different model architectures, sizes, and providers.
6.2.1. Evaluation Dataset: Linux Kernel Driver Analysis with static analyzers
Our study utilizes the Android kernel (Linux version 4.14.150, Google Pixel 4XL), which served as the testbed in the original Suture study (zhang_statically_2021, ). For our initial analysis, potentially informing RQ1 regarding baseline performance and the challenges of automated bug detection in this complex environment, we applied prior static analysis tools, Suture and CodeQL-OOB, to the Linux kernel device drivers. The key findings from this analysis provide important context:
  • $\bullet$
    Suture: When applied to the Linux device drivers, Suture initially generated 251 potential bug reports. This raw output translates to a high False Positive (FP) rate of approximately 90%. Suture employed a subsequent semi-automated refinement process, reporting a reviewer-perceived FP rate of 51.23%. However, this figure relies on Suture’s broader definition of a bug, which classified all integer overflows as true positives. Furthermore, during our investigation, we identified 4 additional instances, initially dismissed as FPs by Suture authors’ verification, that were indeed real bugs (by either their standard or ours). This finding highlights potential inconsistencies in large-scale manual verification efforts.
  • $\bullet$
    CodeQL-OOB: We leverage CodeQL (codeql, ) as a complementary static analysis tool beyond Suture. Based on Backhouse et al.’s approach (backhouse_stack_2018, ), we implemented a simple inter-procedural taint tracking analysis (CodeQL-OOB) that traces data flows from ioctl entry points to pointer dereference and copy_from_user (two typical cases of OOB bugs) to detect potential stack overflows. Running CodeQL-OOB on the Linux device drivers, it yields 24 potential bug reports. Our manual analysis confirmed 1 of these as a true positive (consistent with the known stack overflow bug reported in (backhouse_stack_2018, )). This corresponds to an FP rate of 95.8%.
6.2.2. Cost
On average, the cost of running BugLens is about \$0.1 per case, under the latest version of OpenAI o3-mini. Each case takes about a few minutes to complete.
6.3. RQ1: Effectiveness
Table 1. Performance of BugLens on top of Suture and CodeQL-OOB.
Method TP TN FP FN Prec Rec F1
Suture 24 0 227 0 0.10 - -
SutureRP 20 202 25 4 0.44 0.83 0.58
SutureBugLens 24 218 9 0 0.72 1.00 0.84
CodeQL-OOB 1 0 23 0 0.04 - -
CodeQL-OOBBugLens 1 22 2 0 0.33 1.00 0.5
6.3.1. Precision and Recall
Table 1 shows the performance for the evaluated two static analyzers and our post-refinement method, BugLens. The results show BugLens significantly enhances precision of both Suture (0.10) and CodeQL-OOB (0.04). For CodeQL-OOB, BugLens increases precision substantially to 0.33 while not missing any real bugs detected before. For Suture, the precision is increased to 0.72 due to a drastic reduction in false positives (from 227 to 9). This refinement does not miss any exisiting bugs.
Moreover, noting that the semi-automated method in Suture, noted as SutureRP, actually shows a lower recall (0.83) than the BugLens refinement (1.0). This is because after examining the positive results of SutureBugLens, we found 4 cases of real bugs that were incorrectly classified as false positives during the manual inspection process in SutureRP.
6.3.2. New Bugs
As mentioned earlier, we found 4 more cases that are real bugs, which previously classified as false positive by human inspection in Suture. Two bugs are from the sound subsystem, reported to the maintainers, while waiting for their feedback. One of them involves a data constraint that appears to sanitize a tainted value but can be bypassed due to subtle control-flow logic.
The other two bugs are from the i2c subsystem. They involve a condition where two tainted values must simultaneously satisfy specific constraints—a case that standard taint analyses typically miss due to their focus on single-source propagation.
We have reported all cases following responsible disclosure. Full technical details will be made available after the issues are resolved.
6.3.3. Analysis of FPs
Despite the general effectiveness of BugLens, it shows 10 FPs. Upon careful examination of these cases, we attribute the inaccuracies to several distinct factors:
  • $\bullet$
    Static Analysis Fundamental Limitations (5 cases): False positives arising from inherent limitations in the underlying static analyzers that BugLens is not designed to address. These include imprecisions in taint tracking through complex data structures, incorrect indirect call resolution, etc. BugLens intentionally operates on the dataflow provided by the static analyzers rather than attempting to verify the accuracy of this information itself.
  • $\bullet$
    Environment and Language Understanding (4 cases): Imprecision resulting from the LLM’s incomplete grasp of C language semantics, hardware-level interactions, and kernel-specific programming patterns.
  • $\bullet$
    Internal Modeling Errors (1 case): Inaccuracy originating from a faulty prediction by an internal analysis component (LLM).
6.3.4. Analysis of FNs
Despite the number of FN shown in the table 1 is 0, BugLens still produced several FNs but gets mitigated by majority voting. Specifically, we observed that the FNs were concentrated in the Constraint Assessor (ConA) component, and the SecIA component typically does not generate FNs due to the arbitrary control hypothesis (AC-Hypo), which will be discussed in §6.4.2.
We identified two primary reasons for these FNs:
  • $\bullet$
    Overlooked Complex Conditions: The LLM sometimes failed to recognize complex conditions of the program and therefore summarize the pre-/postcondition incorrectly.
  • $\bullet$
    Misinterpreting Validation vs. Sanitization: LLMs occasionally misclassified sanitization as validations, thereby missing vulnerabilities.
These specific failure patterns observed within ConA were the direct motivation for designing the Structured Analysis Guidance (SAG). The SAG mechanism enhances the prompts used specifically within the ConA stage, providing targeted instructions aimed at guiding the LLM to avoid these identified pitfalls (e.g., explicitly probing for bypass logic, carefully differentiating data constraint types).
The positive impact of SAG in mitigating these FNs is empirically demonstrated in RQ2 (§6.4). As shown in Table 2, the Full Design configuration (using ConA with SAG prompts) consistently yields fewer FNs.
Nevertheless, the impact of SAG is not absolute. The ability of a language model to follow such complex instructions can vary significantly, particularly across different models. Our results in RQ2/RQ3 (Table 2) indicate that some models could still cause some residual FNs even with the SAG.
Takeaway 1. BugLens can effectively post-refine the results of exisiting static analyzers, hugely improving the precision, and can even find missed bugs.
Table 2. Bug Analysis Performance Comparison Across LLMs and Design Approaches (Total Cases=120, Real Bugs=22)
Model Full Design w/o SAG Simple Prompt
FN FP F1 FN FP F1 FN FP F1
[…]
Gemini 2.5 Pro 12 3 0.57 14 4 0.47 6 24 0.52
Claude 3.7 Sonnet 13 2 0.54 17 2 0.34 1 51 0.44
[…]
Table 3. Performance of SecIA, with and without the Aribitary Control Hypothesis (AC-Hypo).
Model w/o AC-Hypo w/ AC-Hypo
FP FN Prec Rec FP FN Prec Rec
OpenAI o3-mini 13 5 0.57 0.77 38 0 0.37 1.0
OpenAI o1 8 4 0.69 0.82 39 0 0.36 1.0
OpenAI GPT-4.1 15 2 0.57 0.91 36 1 0.69 0.95
Gemini 2.5 Pro 60 3 0.24 0.86 73 0 0.23 1.0
Claude 3.7 Sonnet 20 2 0.50 0.91 31 0 0.42 1.0
DeepSeek R1 25 12 0.29 0.45 71 4 0.18 0.82
6.4. RQ2: Component Contribution
To address RQ2, we conduct an incremental analysis. This study evaluates the contribution of the Security Impact Assessor (SecIA), the subsequent Constraint Assessor (ConA), and the specialized Structured Analysis Guidance (SAG) design used within ConA, by comparing performance across progressively enhanced configurations of BugLens.
We assess the performance under the following configurations:
  • $\bullet$
    Baseline: This configuration employs the simple prompt design, based on related works (wen_automatically_2024, ; li_iris_2025, ). It establishes the baseline performance relying primarily on the LLM’s inherent capabilities with minimal guidance. The baseline design provides a starting point for comparison.
  • $\bullet$
    + SecIA: Adds the SecIA component to the Baseline. Purpose: Comparing this to the Baseline isolates the contribution of the SecIA stage. Detailed metrics for SecIA’s filtering rate and soundness are in Table 3.
  • $\bullet$
    + SecIA + ConA (w/o SAG): Adds the ConA component to the ”+ SecIA” configuration, utilizing a simpler prompt design for constraint checking (i.e., without SAG). Comparing this to “+ SecIA” isolates the contribution of adding the constraint assesses step itself.
  • $\bullet$
    Full Design (+ SecIA + ConA + SAG): This configuration enhances the ConA component from the previous step by incorporating the specialized SAG design. This represents the complete BugLens system. Comparing this to previous configurations emphasize the contribution of the SAG.
For this component analysis (RQ2) and the subsequent model versatility analysis (RQ3), we focus our evaluation on a dataset derived from Linux kernel analysis, specifically targeting the sound module. This module was selected because the original Suture study identified it as containing a high density of true positive vulnerabilities (22 out of 24 known bugs), providing a rich testbed for assessing bug detection capabilities. The dataset consists of 120 cases, with 22 known bugs (positives) and 98 non-bug cases (negatives).
We evaluate the performance of these components for diverse LLMs, including OpenAI’s o3-mini, o1, GPT-4.1, Gemini 2.5 Pro, Claude 3.7 Sonnet, and DeepSeek R1. The overall performance results for these configurations are summarized in Table 2, while Table 3 provides the detailed breakdown specifically for the SecIA component’s effectiveness and filtering metrics.
6.4.1. Baseline
Our experimental results clearly demonstrate the significant contribution of our proposed multi-phase workflow and its components compared to a baseline approach. As shown in Table 2, despite some models like Claude 3.7 Sonnet (FN=1) and GPT-4.1 (FN=3) showed low False Negatives, potentially reflecting their raw analytical power, this came at the cost of high False Positives (FP=51 and FP=23, respectively), rendering this simple design ineffective for practical use. The F1 scores for the baseline were generally low across models. This direct prompting approach demonstrated worse performance when compared to other BugLens configurations.
6.4.2. Security Impact Assessor (SecIA)
As Table 3 shows, our Arbitrary Control Hypothesis (AC-Hypo) enhances recall across all models. Without AC-Hypo, the models exhibit noticeable False Negatives, ranging from 2 to 12 FN cases across tested models. After applying AC-Hypo, the FN rate decreases to zero for all models except DeepSeek R1 and GPT-4.1, which maintains a low FN rate (4 and 1, respectively), achieving a high recall of 0.82 and 0.95.
Meanwhile, SecIA demonstrates strong effectiveness as a security vulnerability filter. Taking OpenAI’s o3-mini as an example, among a total of 98 negative cases, SecIA successfully filters out 60 cases (TN). This indicates that SecIA not only has a high recall rate, but it is also effective, substantially improving analysis efficiency.
6.4.3. Constraint Assessor (ConA) without Structured Analysis Guidance (SAG)
While this multi-phase workflow (i.e., SecIA + ConA) significantly reduces the high volume of FPs seen in the Baseline; for instance, Claude 3.7 Sonnet’s FPs dropped from 51 to 2, and Gemini 1.5 Pro’s from 24 to 4. It also leads to a significant increase in False Negatives (FNs) for Gemini 2.5 Pro (FN=14), Claude 3.7 Sonnet and DeepSeek R1 (FN=17), and OpenAI o1 (FN=10) in the ‘w/o SAG’. This supports our hypothesis (§LABEL:subsubsec:naive-llm) that providing constraints, while helpful for pruning obvious non-bugs, can encourage LLMs to become overly confident. Once patterns suggesting data validity are identified, the LLM may default to classifying the issue as “not a bug,” reflecting a potential statistical bias towards common safe patterns rather than performing nuanced reasoning about subtle flaws or bypass conditions.
6.4.4. Structured Analysis Guidance (SAG)
Comparing the Full Design (using SAG within ConA) to the w/o SAG configuration in Table 2 demonstrates SAG’s effectiveness. Introducing SAG leads to a substantial reduction in FNs across all tested models. Consequently, the overall F1 score sees a marked improvement with SAG (e.g., improving from 0.67 to 0.94 for o3-mini and 0.34 to 0.54 for Claude). This indicates that SAG successfully guides the LLM within ConA to overcome the previously observed overconfidence, achieving a better balance between FP reduction and FN mitigation.
Takeaway 2. The design components of BugLens enables effective LLM bug analysis by significantly reducing both FP and FN compared to baseline prompting.
6.5. RQ3: Model Versatility
The results shown in Table 2 affirm that BugLens is a general LLM-based technique applicable across different models, consistently improving upon baseline performance. However, the degree of success highlights variations in how different LLMs interact with complex instructions and structured reasoning processes.
As noted in RQ2, the baseline performance offers a glimpse into the models’ raw capabilities, somewhat correlating with general LLM benchmarks where Gemini 2.5 Pro and Claude 3.7 Sonnet are often considered leaders (vellum2025leaderboard, ). However, this raw capability did not directly translate to superior performance within our structured task without significant guidance.
When employing the ‘Full Design’ of BugLens, we observed distinct differences in instruction-following adherence. The OpenAI models, o1, GPT-4.1 (F1=0.81), and particularly our core model o3-mini (F1=0.94), demonstrated excellent alignment with the workflow’s intent.
Conversely, while the ‘Full Design’ significantly improved the F1 scores for Gemini 2.5 Pro (0.57) and Claude 3.7 Sonnet (0.54) compared to their baseline or ‘w/o sag’ results by drastically cutting down FPs, they still struggled with relatively high false negatives (FN=12 and FN=13, respectively). This suggests that even with the SAG, these powerful models may face challenges in precisely balancing the various analytical steps or interpreting the nuanced instructions within our workflow, possibly still exhibiting a degree of the over-confidence (for “sanity check”) that SAG could not fully overcome in their case. DeepSeek R1 (F1=0.77) showed a strong, balanced improvement, landing between the GPT models and the Gemini/Claude in terms of final performance with the full design. This demonstrates the generality of the BugLens to open-sourced models.
In summary, while our approach is broadly applicable, its optimal performance depends on the LLM’s ability to robustly follow complex, multi-step instructions, with models like OpenAI’s o3-mini currently showing the strongest capability in this specific structured bug analysis task.
Takeaway 3. BugLens shows broad applicability and yields promising results across diverse LLMs, including the open-source DeepSeek R1. OpenAI’s o3-mini currently gets the best result.
6.6. Case Study: Data Structure Traversal
Linux kernel code often uses pointers to traverse data structures. For example, the following code walks through a linked list list using a marco list_for_each_entry:
In this code, the loop goes through each element in the list and checks if any of them match the input id, which comes from the user. This kind of loop is hard for static analysis tools to understand unless they are specially designed to handle linked lists. These tools might wrongly report a warning, saying that user input affects when the loop ends. But in reality, the loop always stops at the end of the list, not because of the user input.
The user input only decides which element gets picked, not how long the loop runs. Our method, BugLens, avoids this false warning by using the LLM’s deeper understanding of how data structures like linked lists work.
6.6.1. A Bypassable Condition
There’s also a tricky part in this code. Before the list_for_each_entry loop even runs, there is a condition that checks if the id->numid is not zero. If it is not zero, the function will return the result of snd_ctl_find_numid(card, id->numid).
This means the checks inside the loop for id->index might not happen at all. So even the loop contains conditions seem to validate id->index, those checks is not always are only performed if id->numid is zero. Otherwise, the function skip the loop entierly.
This cases directly causes false negatives in BugLens, especially without the SAG mechanism. Limited by space, we describe this case in detail in the appendix §A.1.
7. Limitation & Discussion
Dependency on LLMs. The performance of BugLens is influenced by the capabilities of the underlying LLMs used for reasoning. While this dependency could potentially impact external validity, our experiments demonstrate BugLens’s robustness across different models. We have evaluated BugLens with multiple LLMs, including both closed-source and open-source models like Deepseek R1, which achieves approximately 80% of the performance of top-performing models. These results confirm that BugLens’s approach generalizes well across different models, though performance variations exist.
Limited Checkers and Bug Scope. Our evaluation uses only taint-style checkers—those inherited from and Suture (and Dr.checker), plus one CodeQL port—focused on memory-safety and DoS bugs in the Linux kernel. This narrow setup threatens external validity, and the performance we report may not carry over to other bugs (e.g., data races) or analysis frameworks. Although the extra CodeQL checker suggests BugLens can transfer across tooling, a broader study is needed for general applicability.
Towards More Sound Analysis. The soundness of current implementation could be improved through two directions: (1) Adding symbolic verification (bhatia_verified_2024, ; cai_automated_2025, ) to validate the LLM’s reasoning and refine outputs based on formal methods (2) Implementing a hybrid architecture where the LLM performs initial code slicing while symbolic execution handles constraint analysis, combining the LLM’s contextual understanding with provably sound formal reasoning.
Intergation with More Static Analysis Tools. The CodeQL-OOB is an interesting static analysis tool that is based on the lightweight static analysis framework, CodeQL. Considering that BugLens can effectively improve the precision of existing static analysis tools, we could implement more corsare-grained (with low precision) static analysis tools and integrate them into BugLens. Compared to heavy static analysis tools such as suture, an imprecise static analysis could be much easier to implement and maintain.
8. Related work
LLM for Program Analyses & Bug Detection LLMs have been widely applied to program analysis tasks for bug detection. IRIS (li_iris_2025, ) combines LLMs’ contextual understanding with CodeQL queries to enhance taint analysis. LLM4SA (wen_automatically_2024, ) leverages LLMs to refine static analysis results. Their approach serves as a baseline for BugLens (referred as simple prompt). LLMDFA and LLMSAN (DBLP:conf/emnlp/WangZSX024, ; DBLP:conf/nips/WangZSXX024, ) use LLMs to perform data flow analysis and bug detection. LLift (llift, ) focuses on use-before-intialization bugs and prompts LLMs to identify and summarize possible initializers. Focusing on real-world problems, BugLens upgrades the scope to a general taint-style bug detection cross multiple functions.
LLMs for program analysis. LLMs have been widely applied to program analysis tasks, including static semantics analysis (DBLP:conf/nips/WangZSXX024, ; DBLP:journals/corr/abs-2501-18160, ), indirect call resolution (cheng_semantic-enhanced_2024, ), and various inference tasks in program verification and synthesis  (chow_pyty_2024, ; le-cong_can_2025, ; pei_can_2023, ; misu_towards_2024, ; cai_automated_2025, ; li_iris_2025, ). Similarly, BugLens leverages LLMs to reason about security impacts, utilizing their contextual understanding of code semantics to enhance the precision of program analysis results.
Reasoning for LLMs. Despite their success on many tasks, the ability of LLMs to reason about code semantics and behaviors remains an active area of research (ding2024semcoder, ; zelikman2022star, ; li2025codei, ; ni2024next, ). Recent studies have shown that LLMs are still far from performing reliable code reasoning, and their predictions are thus fragile and susceptible to superficial changes in input (pei2023exploiting, ; steenhoek_err_2025, ; hochlehnert2025sober, ). This fragility is often attributed to the learned models taking “shortcuts” based on superficial patterns in training data rather than robust, generalizable reasoning strategies (mccoy_embers_2024, ; yefet2020adversarial, ; gao2023discrete, ; yang2022natural, ; bielik2020adversarial, ). BugLens mitigates this problem with a similar spirit to existing works on boosting the LLMs’ reasoning by constraining their reasoning space with structural and symbolic procedures (li2025structured, ; chen2022program, ; li2023chain, ; chen2024steering, ; yao2023tree, ).
9. Conclusion
This paper introduces BugLens, an innovative post-refinement framework that integrates Large Language Models (LLMs) with static analysis. By employing Security Impact Assessor (SecIA), Constraint Assessor (ConA), and Structured Analysis Guidance (SAG) to guide LLMs through the reasoning process, BugLens significantly enhances the precision of initial static analysis findings without sacrificing scalability. Our evaluation demonstrates that BugLens dramatically reduces false positives in Linux kernel Analysis, minimizes manual inspection effort, and uncovers previously ignored vulnerabilities, highlighting the promise of guided LLMs in making automated bug detection more practical and effective.
Appendix A Case Study of BugLens
A.1. Data Structure Traversals & Bypass Conditions
In §6.6, we have shown a case of data structure traversal:
We now consider a caller of our example snd_ctl_find_id. The function snd_ctl_elem_write takes a control parameter, which is also tainted user input. It then calls snd_ctl_find_id to find the corresponding kctl object. The put function of kctl is then called with the tainted control parameter.
A.1.1. Step 2: Constraint Collection
Suppose the put function is a sink or critical operations that could lead to a vulnerability (and there are indeed many such cases, for different instances of kctl).
When Constraint Assessor (ConA) analyzes this code, its second step is to collect the data constraints of for the tainted input control (actually, the tainted input is control->id, and specific field depends on how the exactly put uses).
Natrually, it will collect the following constraints: snd_ctl_find_id(...) cannot return NULL. and the kctl object must be valid.
A.1.2. Step 3: Constraint Effect Summarization
After collecting the constraints, the next step is to summarize the effect of the constraints, so we need to go back to the snd_ctl_find_id function and analyze the constraints of control->id.
Here’s the most triky part, suppose we track the taint data control->id.index and by looking at the definition of snd_ctl_find_id, we can see that:
Naively, LLM may think these two conditions ensures the control->id.index is within a strict range, between kctl->id.index and kctl->id.index + kctl->count Considering that these kctl objects are predefined and maintained in the kernel, we might think our tainted input is used with an effective and restricted range.
A Closer Look: However, this is not the case. The key here is before the list_for_each_entry loop:
Before the loop, the function checks if id->numid is not zero, and if so, it calls snd_ctl_find_numid to find the corresponding kctl object, and returns it directly. This means that the list_for_each_entry loop, and inside checks can be bypassed if the id->numid is not zero.
The function snd_ctl_find_numid is defined as:
This function also traverses the card->controls list, but it uses kctl->id.numid as the key to find the corresponding kctl object. The numid is a unique identifier for each kctl object, and it is not directly related to the control->id.index field.
Finally, the current effect summarization for the snd_ctl_find_id function should be like (not shown completely):
  1. (1)
    Precondition: control->id.numid != 0,
    Postcondition: kctl != NULL && kctl->id.numid <= control->id.numid < kctl->id.numid+kctl->count
  2. (2)
    Precondition: control->id.numid == 0
    Postcondition: kctl != NULL && kctl->id.index <= control->id.index < kctl->id.index+kctl->count
The overlook of this bypass is the main reason of the false negatives for models.
Appendix B Outline of LLM Prompt
We provide the outline of the LLM prompt for each task in BugLens. For simplified, we remove all examples (that used in few-shot in-context learning).
The args are the arguments for each case running, and the callback is the callback function for each prompt for the agentic design PKA.
B.1. Prompt for Inferring Variable Names
I have a static analysis tool that tracks tainted user input in Linux kernel drivers.
Since the analysis uses LLVM IR, I need help identifying the corresponding source-level variables.
For each case:
1. Tainted values are either local variables loaded from globals/parameters, or direct function parameters
2. Tainted values are always numeric
3. Be aware of name redundancy, especially in global variables and their fields
4. Identify specific struct field names when applicable
5. For propagation, identify the first local variable receiving the taint
Analyze line by line:
The bug detector is: {}
source code:
Instructions:
line no (in source code): {}
args:
- get_bug_detector
- get_function_first_part
- get_insts_from_ctx
- get_source_line_set
B.2. Prompt for SecIA
I have a static analysis tool for Linux kernel drivers that produces many false positives when detecting:
1. Tainted Arithmetic Operations (integer overflows)
2. Tainted Loop Bound Conditions (infinite loops, unexpected iterations)
3. Tainted Pointer Dereferences (arbitrary memory access)
4. Buffer Overflow (out-of-bounds access)
5. Tainted length in copy_from_user (especially stack overflow)
Your Task:
1. Analyze the provided code snippet flagged by our tool
2. Assume:
- Attackers cannot control the kernel
- Tainted variables can be set to any value within their type range
- All existing checks can be bypassed
- Ignore any security checks in the code
3. Determine if the code represents:
- Potential Bug: If the tainted variable can cause infinite loops, very large loops, or memory bugs in the current context
- Normal Code: If the tainted variable usage doesnt lead to security issues within our scope
Provide a step-by-step explanation of your reasoning and classification.
tainted_varaible: {}
bug detector: {}
source code
args:
- get_tainted_value
- get_bug_detector
- get_function
B.2.1. Schema-constrained summarization
Summarize our discussion, and respond with a <bug_eval> tag indicating whether the tainted variable can lead to vulnerabilities:
currently we only consider (1) infinite/very-large loop, (2) out-of-bound access/buffer overflow/arbitrary memory access;
for other types of bugs, you can say "normal code" or "not_a_bug"
<bug_eval>
<tainted_var>tainted_var</tainted_var>
<vulns>
<vuln>
<type>out_of_bound_access</type>
<desc>Brief description of how the vulnerability occurs</desc>
</vuln>
</vulns>
</bug_eval>
For no vulnerability: <bug_eval>not_a_bug</bug_eval>
For potential vulnerability: <bug_eval>potential_bug</bug_eval>
If uncertain: <bug_eval>uncertain</bug_eval>
B.3. Prompt for ConA
This part we provide the prompt for the Constraint Assessor (ConA) task. These prompts already contain the design of SAG. Additionally, we also provide a list of functionalities of the agent PKA inside prompts, represented as {AGENT PROMPTS HERE}.
B.3.1. Step 1: Reachability Analysis
The following is the prompt for the step 1 of ConA task. (see §4.2.1 for details)
Identify the sink (the last line of the provided function context) and determine its preconditions.
Focus only on two types of preconditions:
1. Direct checks (dominate conditions) - Conditions that directly control if the sink executes
Example: if (flag) { sink(tainted_var); }‘
2. Early returns/bypasses (guard conditions) - Conditions that cause early returns or jumps that bypass the sink
Example: if (tainted_var > 100) return; sink(tainted_var);‘
Here, "tainted_var <= 100" is the precondition to reach the sink.
Ignore conditions that dont directly impact sink reachability.
{AGENT PROMPTS HERE}
sink variable
sink context (the full context of the function):
Summarize in this format:
‘‘‘xml
<sink_precondi>
<precondi>
<type> dominate_condition </type>
<condition> flag </condition>
<dominated_sink> if(flag) sink(tainted_var) </dominated_sink>
</precondi>
<precondi>
<type> guard_condition </type>
<condition> tainted_var <= 100 </condition>
<guard_bypass> if (tainted_var > 100) return/goto invalid_label; </guard_bypass>
</precondi>
</sink_precondi>
Note: Multiple preconditions combine with "AND" logic.
args:
- get_tainted_value
- get_function
callback:
- need_struct_def
- need_global_var_def
B.3.2. Step 2: Constraint Collection
The following is the step 2 of the prompt for the ConA task. (see §4.2.2 for details)
Help find range constraints for the tainted variable across the call chain.
tainted_variable: {}
current callchain: {}
context of the sink:
Focus on finding all possible value range constraints for the tainted variable, without analyzing their effectiveness.
Types of constraints to look for:
1. Validation: conditions that reject invalid ranges (e.g., if (tainted_var < 0) return -EINVAL;)
2. Sanitization: corrections applied to the value (e.g., tainted_var = min(tainted_var, 100);)
3. Type constraints: implicit limits from variable types (e.g., uint8_t limits to [0,255])
Important notes:
- Validations are transferable through operations (e.g., constraints on var = tainted_var + 1 apply to tainted_var)
- Sanitizations are not transferable
- Base your analysis only on the provided code, not prior knowledge
- Track the variable across different names in different contexts
{AGENT PROMPTS HERE}
Summarize your findings in this format:
<range_constraints>
<constraint>
<type>validation|sanitization|type_constraint</type>
<handler_func>function_name</handler_func>
<context>relevant_code_snippet</context>
</constraint>
</range_constraints>
args:
- get_tainted_value
- get_call_chain
- get_function_first_part
callback:
- need_func_def
- need_caller
- need_struct_def
- need_global_var_def
B.3.3. Step 3: Constraint Effect Summarization
The following is the prompt of the step 3 of the ConA task. (see §4.2.3 for details)
Act as a program verifier/symbolic execution engine to infer precondition and postcondition pairs for constraints in code. Use these simple rules:
- The **precondition** is the path condition required to reach the constraint.
- The **postcondition** is the effect on the tainted variable (often shown as its valid range).
Consider all branches and early return "bypass" cases. For instance, given this sample function:
‘‘‘c
void check_tainted_value(int config, int other_config, int x){
if (other_config == CHECK_SKIP)
return;
if (config == 1)
assert(x > 0 && x < 100);
else if (config == 2)
assert(x > 0);
You must extract the following [precondition, postcondition] pairs:
1. Bypass Case
Precondition: other_config == CHECK_SKIP
Postcondition: x in (-inf, +inf)
2. Config 1 Case
Precondition: other_config != CHECK_SKIP && config == 1
Postcondition: x in (0, 100)
3. Config 2 Case
Precondition: other_config != CHECK_SKIP && config == 2
Postcondition: x in (0, +inf)
4. Default Case
Precondition: other_config != CHECK_SKIP && config != 1 && config != 2
Postcondition: x in (-inf, +inf)
{AGENT PROMPTS HERE}
Finally, output your analysis as XML in the following format:
‘‘‘xml
<range_constraint>
<type>validation</type>
<handler_func>check_tainted_value</handler_func>
<condition_pairs>
<pair>
<precondi>other_config == CHECK_SKIP</precondi>
<postcondi>x in (-inf, +inf)</postcondi>
<context> if (other_config == CHECK_SKIP) return; </context>
</pair>
<pair>
<precondi>other_config != CHECK_SKIP && config != 1 && config != 2</precondi>
<postcondi>x in (-inf, +inf)</postcondi>
<context> induced by the other branches </context>
</pair>
</condition_pairs>
</range_constraint>
callback:
- need_func_def
- need_struct_def
- need_global_var_def
- need_caller
B.3.4. Step 4: Final Vulnerability Evaluation
The following is the prompt of the step 4 of the ConA task. (see §4.2.4 for details)
Evaluate if this bug is eliminated, not exploitable, or still vulnerable.
Extract concrete values for "size of" or "length of" in constraints and bug conditions.
Analyze whether the bug condition can be satisfied considering all constraints:
- Disregard constraints with postconditions that dont limit the tainted variable
- Disregard constraints for unrelated sinks
- For kernel-controlled conditions, determine reachability based on your knowledge
- If a precondition contains user-controlled variables, assume users can bypass it
A bug is "eliminated" only if:
- The postcondition restricts the tainted variable to a safe range that makes the bug impossible
- The precondition is always satisfied when the sinks precondition is true
If the postcondition isnt strong enough, assess exploitability assuming attackers can set any value.
{AGENT PROMPTS HERE}
callback:
- need_func_def
- need_struct_def
- need_global_var_def
- need_caller
B.3.5. Schema-constrained summarization
The following is the prompt for summarization of the ConA task, which is used to extract the final result of the vulnerability evaluation within a <final_res>.
Now Lets summarize our discussion, and respond in a <final_res> tag with the following format:
"still_a_bug", "eliminated", "likely_safe", "likely_unsafe", "not_exploitable" or "uncertain" within a <final_res> tag, e.g., <final_res>still_a_bug</final_res>
B.4. Prompt for PKA
The following is the exact prompt that shown as AGENT PROMPTS HERE in the previous parts.
1========================================
2First of all, you dont need to complete the task in your initial response. You can always ask for more information.
3
4When you need additional details, use the following format. In this case, dont reach a conclusion immediately - instead, request the information you need to perform a thorough analysis once you receive my response.
5--- request 1: ask for the function definition ---
6You could ask me for the definition of the function. in this case, you could respond with the following:
7<requests>
8 <request>
9 <name>need_func_def</name>
10 <args>
11 <arg>func_1</arg>
12 <arg>func_2</arg>
13 </args>
14 </request>
15</requests>
16--- request 2: ask for the struct definition ---
17You could ask me for the definition of the structure. in this case, you could respond with the following:
18<requests>
19 <request>
20 <name>need_struct_def</name>
21 <args>
22 <arg>struct_name_1</arg>
23 <arg>struct_name_2</arg>
24 </args>
25 </request>
26</requests>
27--- request 3: ask for the caller of the current function ---
28You could ask me for the caller for the current function. (Note: you can only request one caller at a time)
29 in this case, you could respond with the following:
30 <requests>
31 <request>
32 <name>need_caller</name>
33 <args>
34 <arg>current_function_name</arg>
35 </args>
36 </request>
37</requests>
38--- request 4: ask for the definition of global variables ---
39You could ask me for the definition of global variables. In this case, you could respond with the following:
40<requests>
41 <request>
42 <name>need_global_var_def</name>
43 <args>
44 <arg>global_var_1</arg>
45 <arg>global_var_2</arg>
46 </args>
47 </request>
48</requests>
49========================================